### v1

Abstract
For a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers, we consider curve and divisorial valuations on the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables with the coeficients in $\mathbb{K}$. We compute the semigroup Poincaré series and the classical Poincaré series of a valuation.
1 Introduction
A (discrete, rank one) valuation on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of germs of holomorphic functions in two variables is a map $\nu:{\mathcal{O}}_{\mathbb{C}^{2},0}\to\mathbb{Z}_{\geq 0}\cup\{+\infty\}$ such that
  1. 1)
    $\nu(0)=+\infty$;
  2. 2)
    $\nu(\lambda f)=\nu(f)$ for $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, $\lambda\in\mathbb{C}^{*}:=\mathbb{C}\setminus\{0\}$;
  3. 3)
    $\nu(f_{1}+f_{2})\geq\min\{\nu(f_{1}),\nu(f_{2})\}$ for $f_{1},f_{2}\in{\mathcal{O}}_{\mathbb{C}^{2},0}$;
  4. 4)
    $\nu(f_{1}f_{2})=\nu(f_{1})+\nu(f_{2})$ for $f_{1},f_{2}\in{\mathcal{O}}_{\mathbb{C}^{2},0}$.
There are essentially two classes of valuations on ${\mathcal{O}}_{\mathbb{C}^{2},0}$: curve valuations and divisorial ones (see, e. g., [11]).
The notion of the Poincaré series of a collection $\{\nu_{i}\}$, $i=1,\ldots,r$, of valuations was introduced in [3]. For $r=1$ it reduces to
$P_{\nu}(t)=\sum_{v=0}^{\infty}\dim_{\mathbb{C}}(J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1))\cdot t^{v}\in\mathbb{Z}[[t]]\;,$ (1)
where $J^{\mathbb{C}}(v):=\{f\in{\mathcal{O}}_{\mathbb{C}^{2},0}:\nu(f)\geq v\}$. The Poincaré series of one curve valuation was known to specialists for many years. Computations can be found, e. g., in [4]. The Poincaré series of one divisorial valuation was computed in [9]. The Poincaré series of a collection of curve valuations was computed in [5]. It was shown that this series coincides with the Alexander polynomial in several variables of the corresponding algebraic link. This implies that the Poincaré series determines the topology of the union of the corresponding curves.
In [7], there were considered valuations on the algebra ${\mathcal{E}}_{\mathbb{R}^{2},0}$ of germs of real analytic functions on $(\mathbb{R}^{2},0)$. (All valuations on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ are restrictions of valuations on ${\mathcal{O}}_{\mathbb{C}^{2},0}$.) There were considered several versions of the Poincaré series (including the one defined in [3]). There was computed the Poincaré series for one curve valuation. In[8], there was computed the Poincaré series of one divisorial valuation in this setting.
2 Curve valuations on functions over a field
Let $\mathbb{K}$ be a subfield of the field $\mathbb{C}$ of complex numbers and let ${\mathcal{E}}_{\mathbb{K}^{2},0}$ be the algebra of germs of holomorphic functions in two variables with the Taylor coefficients from $\mathbb{K}$. A (discrete, rank one) valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ (defined as above) is the restriction of a valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$: see, e. g., [10].
In the framework of the discussion below, without changes it is possible to consider, instead of the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$, the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables. We shall refer to ${\mathcal{E}}_{\mathbb{K}^{2},0}$ because of some convenience of the language.
Let $(C,0)$ be a plane algebroid branch on $(\mathbb{C}^{2},0)$, that is $C$ (after an appropriate change of the coordinates in $\mathbb{C}^{2}$) is defined by
$x=x(\tau)=\tau^{m},\\ y=y(\tau)=\sum_{i\geq m}c_{i}\tau^{i}\in\mathbb{C}[[\tau]]\,.$ (2)
(We assume that the parametrization (2) is irreducible, i. e., the greatest common divisor of $m$ and of all $i$ such that $c_{i}\neq 0$ is equal to $1$.) The branch $C$ defines a valuation $\nu=\nu_{C}$ on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of holomorphic function germs in $x$ and $y$. For $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, $\nu(f)$ is the degree of the leading term in the series $f(x(\tau),y(\tau))\in\mathbb{C}[[\tau]]$:
$f(x(\tau),y(\tau))=a(f)\cdot\tau^{\nu(f)}+\text{ terms of higher degree}\;,$ (3)
where $a(f)\neq 0$. If $f(x(\tau),y(\tau))\equiv 0$, $\nu(f)):=+\infty$. Let us consider its restriction to the subalgebra ${\mathcal{E}}_{\mathbb{K}^{2},0}\subset{\mathcal{O}}_{\mathbb{C}^{2},0}$. For $v\in\mathbb{Z}_{\geq 0}$, let $J^{\mathbb{K}}(v):=\{f\in{\mathcal{E}}_{\mathbb{K}^{2},0}:\nu(f)\geq v\}$. If all the quotients $J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)$ are finite-dimensional (over $\mathbb{K}$) the definition in [3] reduces to the following one:
Definition 1.
The $\mathbb{K}$-Poincaré series of the branch $C$ is
$P_{C}(t)(=P_{C}^{\mathbb{K}}(t))=\sum_{v=0}^{\infty}\dim_{\mathbb{K}}\left(J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\right)\cdot t^{v}.$ (4)
Example 1.
If $\mathbb{K}=\mathbb{R}$, $P_{C}(t)$ is the classical Poincaré series of the valuation $\nu$ computed in [7].
Remark 1.
If $\mathbb{K}=\mathbb{C}$, all the coefficients of the series (4) are equal to 00 or $1$. This is not the case, in general, for a proper subfield $\mathbb{K}\subset\mathbb{C}$. In particular, if $\mathbb{K}=\mathbb{R}$ (the field of real numbers), all the coefficients are equal to 00, $1$, or $2$ ([7]).
Let $S^{\mathbb{K}}_{C}:=\{\nu(f):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}\}\subset\mathbb{Z}_{\geq 0}$ be the semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
Definition 2.
The $\mathbb{K}$-semigroup Poincaré series of the valuation $\nu$ is $P^{S}_{C}(t)=\sum_{v\in S^{\mathbb{K}}_{C}}t^{v}$.
Example 2.
For $\mathbb{K}=\mathbb{R}$ this is the semigroup Poincaré series considered (and computed) in [7].
Here we compute the $\mathbb{K}$-semigroup Poincaré series and the $\mathbb{K}$-Poincaré series for one curve or divisorial valuation.
3 The resolution process of a branch
Let $G$ be the Galois group of the extention $\mathbb{C}/\mathbb{K}$. i. e. the group of automorphisms of $\mathbb{C}$ which are trivial on $\mathbb{K}$. Let us consider the minimal $G$-invariant resolution process of the branch $C$ (by a sequence of blow-ups). This means the following. When we blow-up the origin in $\mathbb{C}^{2}$, we have got one of the following three possibilities.
  1. 1)
    There are finitely many different branches in the $G$-orbit of the strict transform $\widetilde{C}$ of the branch $C$, all of them are smooth and they intersect the exceptional divisor transversally at pairwise different points. We get (the minimal embedded $G$-invariant) resolution of $C$ and the process is finished.
  2. 2)
    The branches from the $G$-orbit of $\widetilde{C}$ (whose number may be infinite) intersect the exceptional divisor at finite number of points, but either they are not smooth, or do not intersect the exceptional divisor transversally, or (some) different branches from the orbit intersect it at the same point. In this case one has to continue the process.
  3. 3)
    The branches from the $G$-orbit of $\widetilde{C}$ intersect the exceptional divisor at infinitely many different points. In this case the process is terminated.
In Case 2 one has to blow-up (all) the intersection points of the branches from the $G$-orbit of $\widetilde{C}$ with the exceptional divisor (their number being finite). One gets one of the analogues the cases 1–3 again. Etc.
As the result we arrive to one of the following three situations.
  1. I
    There are finitely many branches in the $G$-orbit of $C$ and they are resolved at a certain step. The dual graph of the resolution looks like in Figure 1.
Figure 1: The resolution graph $\Gamma$ in case I (here $s=2$, $g=6$, see below).
The Galois group $G$ acts on the graph $\Gamma$ and the quotient $\check{\Gamma}$ by the action looks like in Figure 2 (and further on).
The graph $\Gamma$ contains a subgraph which geometrically coincides with the dual graph of a usual resolution of the curve $C$: the lower part of $\Gamma$ on Figure 1. (We do not consider this resolution below.) We use “short” notations (by greek letters) for verticies on this subgraph (and also for the corresponding verticies of the graph $\check{\Gamma}$). All other verticies are obtain from them by the action of elements of the Galois group $G$.
In Figure 1, $\sigma_{i}$, $i=0,\ldots,g$, (and also their images under the action of $G$) are the dead ends of the graph ($g$ is the number of the Puiseux pairs of the curve $C$), $\tau_{i}$, $i=1,\ldots,g$, (and also their images) are the rupture points of it, $\rho_{j}$, $j=1,\ldots,s$, ($s$ is equal to $\infty$ in Case II) are the splitting points corresponding to the changes of the isotropy subgroups for the $G$-action on $\Gamma$ (that is the splitting points for usual resolution of different branches from the orbit of $C$, there are finitely many of them in Case I infinitely many in Case II.
Figure 2 shows $\check{\Gamma}$ in the case when the process of the $\mathbb{K}$-resolution of $C$ is not finish when $C$ itself is resolved. Otherwise $\delta_{C}$ coincides with $\tau_{g}$.
Figure 2: The resolution graph $\check{\Gamma}$ in Case I (with arbitrary $g$ and $s$).
  1. II
    The described process goes on without termination. This means that we get an infinite sequence of blow-ups. The corresponding dual graph of the process looks like in Figure 3. One can say that $s$ is equal to infinity in this case.
    Figure 3: The resolution graph $\check{\Gamma}$ in Case II (with $s=\infty$).
  2. III
    The process is terminated at a certain step. This means that at this step the branches from the $G$-orbit of $\widetilde{C}$ intersect each of the new-born components of the exceptional divisor at infinitely many points. The dual graph of the process looks like in Figure 4.
    In the latter case, the valuation coincides with $n\nu$, where $\nu=\nu_{\rho_{s}}$ is the divisorial valuation defined by $E_{\rho_{s}}$ and $n$ is the intersection multiplicity $[E_{\rho_{s}},\widetilde{C}]$ of the strict transform $\widetilde{C}$ of the curve $C$ with the component $E_{\rho_{s}}$.
    Figure 4: The resolution graph $\check{\Gamma}$ in Case III.
Cases I–III take place in the following situations.
  1. I
    All the coefficients in the series (2) are from a finite extension of $\mathbb{K}$. This is equivalent to the condition that the orbit $GC$ has finitely many elements.
  2. II
    The coefficients all together are not from a finite extension of $\mathbb{K}$, but each of them is algebraic over $\mathbb{K}$.
  3. III
    There exists (at least) one coefficient transcendental over $\mathbb{K}$.
Remark 2.
Let $\mathbb{K}\subset\mathbb{C}$ (e. g., $\mathbb{K}=\mathbb{Q}$) and let $F\in\mathbb{K}[[x,y]]$ be irreducible. The local ring of the irreducible curve (over $\mathbb{K}$) defined by $F$ is $R=\mathbb{K}[[x,y]]/(F)$. The integral closure $\overline{R}$ in its quotient field is the valuation ring $R_{\omega}$ of a discrete valuation $\omega$, its residual field $\mathbb{K}^{\prime}$ is a finite algebraic extension of $\mathbb{K}$ and could be realized and therefore identified with a subfield of $\mathbb{C}$. The equation $F$ defines a reduced algebroid curve over the complex field $\mathbb{C}$, each branch corresponds to an irreducible factor of $F$ in $\mathbb{C}[[x,y]]$. Let $C$ be one of these branches. The valuation $\nu=\nu_{C}$ defined by $C$ on $\mathbb{K}[[x,y]]$ coincides with the valuation $\omega$ (seen as a valuation on $\mathbb{K}[[x,y]]$ forgeting the quotient by $F$) and the Poincaré series (over $\mathbb{K}$) defined by $\nu$ is the Poincaré series defined by $\omega$. Note that all the branches are equisingular, the minimal resolution of $\{F=0\}$ (in the complex setting) coincides with the $\mathbb{K}$-resolution of $\nu_{C}$ and all the valuations $\nu_{C}$, for complex branches $C$ of $\{F=0\}$, restrict to the same valuation $\omega$ in $\mathbb{K}[[x,y]]$. Moreover, if $G^{\prime}$ is the Galois group of the extension $\mathbb{K}\subset{\mathbb{L}}$, ${\mathbb{L}}$ the normal closure of $\mathbb{K}^{\prime}$ in $\mathbb{C}$, $G^{\prime}$ is the quotient of $G$ by the isotropy group of the field $\mathbb{L}$. This situation just corresponds to the case I described above.
4 Subgroups of the Galois group, subfields, divisors and curvettes
Further on we shall discuss only Cases I and II. Let $(E_{\sigma}\circ E_{\delta})$, $\sigma,\delta\in\Gamma$, be the intersection matrix of the components $E_{\sigma}$ (infinite one in Case II). The diagonal entry $E_{\sigma}\circ E_{\sigma}$ is the self-intersection number of the component $E_{\sigma}$ on the final modification in Case I and on a modification at a step, after which the points of $E_{\sigma}$ are not blown-up, in Case II. One has $E_{\sigma}\circ E_{\sigma}<0$. The non-diagonal entry $E_{\sigma}\circ E_{\delta}$ is equal to 1 (respectively to 00) if these components intersect (do not intersect respectively).
Let $(m_{\sigma,\delta}):=-(E_{\sigma}\circ E_{\delta})^{-1}$. In Case II one can compute the entry $m_{\sigma,\delta}$ at a sufficiently late step (after which it stabilizes). In other words one can compute the intersection matrix and the inverse of it at each step and use their stabilization.
The number $m_{\sigma,\delta}$ can be interpreted as the intersection number of some germs of curves on $(\mathbb{C}^{2},0)$, so-called curvettes. A curvette correponding to the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a complex-analytic smooth curve $\gamma_{\sigma}$ (in Case II on a modification where $E_{\sigma}$ stabilizes) transversal to the component $E_{\sigma}$ at a smooth point of the exceptional divisor, that is not an intersection point with other components. One can show that $m_{\sigma,\delta}$ is equal to the intersection number $C_{\sigma}\circ C_{\delta}$.
For $\sigma\in\Gamma$, let $m_{\sigma}$ be the value of $\nu_{C}$ on the left-hand side of an equation $f_{\sigma}=0$ defining the curvette $C_{\sigma}$, i. e., the intersection number $C\circ C_{\sigma}$. One can see that, in Case I, $m_{\sigma}$ is equal to $m_{\sigma,\delta_{C}}$ for $\delta_{C}$ such that, on the surface of the resolution, the strict transform of the branch $C$ intersects the component $E_{\delta_{C}}$ and, in Case II, $m_{\sigma}$ is equal to $m_{\sigma,\delta}$ for $\delta$ on the subgraph of the usual resolution of the curve $C$ far enough from the initial component of the modification.
Isotropy subgroups and subfields.
Let $G_{\sigma}$ be the isotropy subgroup of the component $E_{\sigma}$, i. e., the isotropy subgroup of the vertex $\sigma$ in $\Gamma$. The subgroup $G_{\sigma}$ has a finite index in the Galois group $G$. Let $\mathbb{K}_{\sigma}\subset\mathbb{C}$ be the invariant subfield of the group $G_{\sigma}$. One can see that, for $g\in G$, $G_{g\sigma}=gG_{\sigma}g^{-1}$, $\mathbb{K}_{g\sigma}=g\mathbb{K}_{\sigma}$. For all the vertices $\sigma\in\Gamma$ inbetween the initial one $\sigma_{0}=$1 and $\rho_{1}$ ($\rho_{1}$ included) the subgroups $G_{\sigma}$ and the subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G=G_{0}$, $\mathbb{K}_{\sigma}=\mathbb{K}=\mathbb{K}_{0}$. For all vertices $\sigma\in\Gamma$ inbetween the vertex $\rho_{j}$ and $\rho_{j+1}$ ($\rho_{j}$ excluded, $\rho_{j+1}$ included) the isotropy subgroups $G_{\sigma}$ and the corresponding subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G_{j}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{j}$. In Case I, for all the vertices $\sigma\in\Gamma$ betwin $\rho_{s}$ and $\delta_{C}$ one has $G_{\sigma}=G_{s}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{s}$. For vertices from $G$-orbits of the discussed ones, the corresponding subgroups are conjugate to $G_{\sigma}$ and the subfields are obtained from $\mathbb{K}_{\sigma}$ by the shift by an element of the Galois group $G$. One has $G=G_{0}\supset G_{1}\supset\ldots\supset G_{s}$, $\mathbb{K}=\mathbb{K}_{0}\subset\mathbb{K}_{1}\subset\ldots\subset\mathbb{K}_{s}$. One has $[K_{j}:K_{j-1}]=[G_{j-1}:G_{j}]$.
Standard charts.
For computations, we shall use coordinate systems compatible with the actions of the Galois groups. This means the following. There are two standard charts covering the initial component of the resolution process: with the coordinates $u=x,w=y/x$ or $u=y,w=x/y$ respectively. If, on a previously created component one makes the blow-up at the point $(u,w)=(0,a)$ (in a standard chart), one has two standard charts covering the new-born component: with the coordinates $u^{\prime}=u,w^{\prime}=(w-a)/u$ and $u^{\prime}=w-a,w^{\prime}=u/(w-a)$ respectively.
On all the components $E_{\sigma}$ of the exceptional divisor except $E_{\rho_{j}}$ and those from its G-orbit, on each step the blow-up is made at one point with the coordinates $(0,a)$ in a standard chart with $a\in\mathbb{K}_{\sigma}$. In this case it can be convenient to change the coordinate $w$ by $\widetilde{w}=w/(w-a)$ or $\widetilde{w}=(w-a)/w$ so that the coordinates of the intersection points with the components on the geodesic between $\sigma_{0}$ and $\delta_{C}$ become equal to 00 and $\infty$. On the component $E_{\rho_{j}}$, the blow-ups are made at a certain point $(0,a)$ with $a\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ (that is not from $\mathbb{K}_{\rho_{j}}=\mathbb{K}_{j-1}$) and at the points from its $G_{j-1}$-orbit. In this case a change as above is inappropriate.
Special curvettes.
In what follows, we shall use curvettes of special type. In a standard chart covering an affine part of $E_{\sigma}$, let the (smooth and transversal) curve $\gamma_{\sigma}$ be defined over the field $\mathbb{K}_{\sigma}$. In other words, the curve $\gamma_{\sigma}$ is invariant under the action of $G_{\sigma}$. In particular one can take the curve $\{u=\tau$, $w=c=\text{const}\}$ in a standart chart with $c\in\mathbb{K}_{\sigma}$ ($u=0$ is the equation of $E_{\sigma}$ in the chart).
Definition 3.
A $\mathbb{K}_{\sigma}$-curvette at the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a curve $\gamma_{\sigma}$ described above.
Remark 3.
One can see that $C_{\sigma}$ has a parametrization like (2) with all the coefficients from $\mathbb{K}_{\sigma}$. A $\mathbb{K}_{\sigma}$-curvette at the component $E_{\sigma}$ can be defined by an equation $f_{\sigma}=0$ with $f_{\sigma}\in{\mathcal{E}}_{\mathbb{K}_{\sigma}^{2},0}$.
Definition 4.
A $G_{\sigma}$-curvette at the component $E_{\sigma}$ is the union of the curves $gC_{\sigma}$ ($C_{\sigma}$ is a $\mathbb{K}_{\sigma}$-curvette at $E_{\sigma}$) over representatives of the $G_{\sigma}$-classes in $G$.
One can show that a $G_{\sigma}$-curvette at the component $E_{\sigma}$ can be defined by an equation $F_{\sigma}=0$ with $F_{\sigma}\in{\mathcal{E}}_{\mathbb{K}^{2},0}$. In fact one can take $F_{\sigma}=\prod_{[g]\in G/G_{\sigma}}(gf_{\sigma})$, where the product is over representatives of the $G_{\sigma}$-classes in $G$.
Remark 4.
Somewhat similar type of symmetries of the resolution graph was considered in [6]. In that case one had an action of a group on the graph, but not on the underlying space $\mathbb{C}^{2}$ or on the space of functions. The series computed in this paper and therein are different. In fact the setting studied here was one of the motivations for [6].
5 The semigroup Poincaré series of a curve valuation
Let $\nu(=\nu_{C})$ be the valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ defined by a curve $C\subset(\mathbb{C}^{2},0)$ and assume that $\nu_{C}$ is a valuation of the type I or II described in Section 3. Let $P^{S}_{C}(t)=\sum_{m\in S^{\mathbb{K}}_{\nu}}t^{m}$ be the semigroup Poincaré series of $\nu_{C}$ (see Definition 2).
For $\sigma\in\check{\Gamma}$, let $M_{\sigma}$ be the value of $\nu$ on the left-hand side of the equation $F_{\sigma}=0$ definig a $G_{\sigma}$-curvette at the component $E_{\sigma}$. One can see that $M_{\sigma}=\sum_{[g]\in G/G_{\sigma}}m_{g\sigma}$. (Compare with the definition of $M_{\sigma}$ in [7], where $\mathbb{K}=\mathbb{R}$.)
The aim of this section is to prove the following statement.
Theorem 1.
$P^{S}_{C}(t)=\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\;.$
Let $S_{C}$ be the usual semigroup of the complex curve $C$, i. e., $S_{C}=\{\nu(f):f\in{\mathcal{O}}_{\mathbb{C}^{2},0}\}$. The set of multiplicities $\{m_{\sigma_{0}},\ldots,m_{\sigma_{g}}\}$ is the minimal set of generators of $S_{C}$. Let $e_{i}:=\gcd\{m_{\sigma_{0}},\,\ldots,m_{\sigma_{i}}\}$ for $i=0,1,\ldots,g$ and let $N_{i}=e_{i-1}/e_{i}$ for $i=1,2,\ldots,g$. (One has $m_{\tau_{i}}=N_{i}m_{\sigma_{i}}$ for $i=1,2,\ldots,g$: see, e. g., [12] or [2].)
For $\sigma\in\Gamma$, let $\pi_{\sigma}:(X_{\sigma},D_{\sigma})\to(\mathbb{C}^{2},0)$ be the minimal modification of $(\mathbb{C}^{2},0)$ such that $E_{\sigma}\subset D_{\sigma}$. In particular, $E_{\sigma}$ is the last exceptional component appearing in $X_{\sigma}$ and is produced by blowing-up at a point $p_{\sigma}$ of a previous component. For $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, we will denote by $e_{\sigma}(f)$ the multiplicity of the strict transform by $\pi_{\sigma}$ of the curve $\{f=0\}$ at the point $p_{\sigma}$. Notice that $e_{\sigma}(f)$ coincides with the intersection multiplicity of the strict transform of $\{f=0\}$ and $E_{\sigma}$ in the surface $X_{\sigma}$.
For $j=1,\ldots,s$, let $\ell_{j}:=[\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]:=|\frac{G_{j-1}}{G_{j}}|$.
Lemma 1.
Let $\omega\in\check{\Gamma}$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$ and $\tau_{p}\leq\omega<\tau_{{p}+1}$. One has
$M_{\omega}=\sum_{i=0}^{{p}}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
for some non negative integers $k_{i}$ $(0\leq i\leq{p})$ and positive integers $r_{j}$ $(1\leq j\leq d)$.
Proof..
Let us assume that $\tau_{q-1}<\rho_{d}\leq\tau_{q}$ (see Figure 5 below). For $i=0,\ldots,g$, let $M_{i}:=M_{\sigma_{i}}$, $m_{i}:=m_{\sigma_{i}}$ and let $\varphi_{i}=0$ be the equation of a curvette at the component $E_{\sigma_{i}}$.
Let $\omega$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$. Then one has
$M_{\omega}$ $=m_{\omega}+([G_{0}:G_{d}]-[G_{1}:G_{d}])e_{\rho_{1}}(\varphi_{\omega})m_{\rho_{1}}+([G_{1}:G_{d}]-[G_{2}:G_{d}])e_{\rho_{2}}(\varphi_{\omega})m_{\rho_{2}}+$
$\qquad\qquad+\cdots+([G_{d-1}:G_{d}]-1)e_{\rho_{d}}(\varphi_{\omega})m_{\rho_{d}}=$
$=m_{\omega}+\sum_{j=1}^{d}([G_{j-1}:G_{d}]-[G_{j}:G_{d}])e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}=m_{\omega}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$
where $L^{d}_{j}=[G_{j-1}:G_{d}]-[G_{j}:G_{d}]$, $1\leq j\leq d$. In the same way, for $i=q,\ldots,{p}$ one has
$M_{i}=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}\,.$ (5)
Figure 5: The position of the vertices.
Moreover, it is known that $m_{\omega}\in\langle m_{0},\,\ldots,m_{p}\rangle$ and so $m_{\omega}=k_{0}m_{0}+\sum_{i=1}^{p}k_{i}m_{i}$, for some $k_{0}\geq 0$ and $0\leq k_{i}<N_{i}$ for $i\geq 1$ (which are unique with this restrictions). Then one can write
$M_{\omega}=\sum_{0}^{p}k_{i}m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$ (6)
Let us assume that $q<{p}$. For $j\leq d$ one has $e_{\rho_{j}}(\varphi_{\omega})/e_{\tau_{{p}-1}}(\varphi_{\omega})=e_{\rho_{j}}(\varphi_{{p}})/e_{\tau_{{p}-1}}(\varphi_{p})$, moreover $e_{\tau_{{p}-1}}(\varphi_{p})=1$, so $e_{\rho_{j}}(\varphi_{\omega})=e_{\tau_{{p}-1}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})$. It is known that $e_{i}=e_{\tau_{i}}(\varphi_{C})$, $i\leq g$ and $\varphi_{C}$ is an equation of the complex curve $C$. As a consequence, for $i\leq p$, one has $N_{i}=e_{i-1}/e_{i}=e_{\tau_{i-1}}(\varphi_{\omega})/e_{\tau_{i}}(\varphi_{\omega})$ and so
$e_{\rho_{j}}(\varphi_{\omega})=N_{p}e_{\tau_{p}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})\;.$
The same computations applied to $\varphi_{i}$ (instead of $\varphi_{\omega}$) and $\varphi_{i-1}$ (for $\tau_{i}>\tau_{i-1}>\rho_{j}$) show that $e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\tau_{i-1}}(\varphi_{i})e_{\rho_{j}}(\varphi_{i-1})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$.
Using this equality in the rigth part of the equation (6) and also the expression of $M_{i}$ above, one gets:
$\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$ $=N_{p}e_{\tau_{p}}(\varphi_{\omega})\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}=$
$=k_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}+k^{\prime}_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}$
with $k^{\prime}_{p}=N_{p}e_{\tau_{p}}(\varphi_{\omega})-k_{p}\geq 1$, and so
$M_{\omega}=\sum_{0}^{{p}-1}k_{i}m_{i}+k_{p}M_{p}+k^{\prime}_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}\;.$
In the same way, using repeatedly that $e_{\rho_{i}}(\varphi_{j})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$ for $i\geq q$ one gets:
$M_{\omega}=(\sum_{0}^{q-1}k_{i}m_{i})+\sum_{i=q}^{p}k_{i}M_{i}+r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}$ (7)
for some integer $r_{d}\geq 1$. Now, using that $L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}$ for $j\leq d-1$ in the last summand of the above equality (7), one has
$r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=r_{d}[G_{d-1}:G_{d}]\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}+r_{d}L^{d}_{d}e_{\rho_{d}}(\varphi_{q})m_{\rho_{d}}\;.$
Using that $L^{d}_{d}=[G_{d-1}:G_{d}]-1$, $e_{\rho_{j}}(\varphi_{q})=e_{\rho_{j}}(\varphi_{\rho_{d}})$, $e_{\rho_{d}}(\varphi_{\rho_{d}})=1$ and the corresponding expression for $M_{\rho_{d}}$ in the same terms as at the beginning of the proof for $M_{\omega}$, the last equality can be written as:
$r_{d}([G_{d-1}:G_{d}]-1)\left(\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{\rho_{d}})m_{\rho_{j}}+m_{\rho_{d}}\right)+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=$
$=r_{d}([G_{d-1}:G_{d}]-1)M_{\rho_{d}}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\;.$
Thus, taking into account that $\ell_{j}=[G_{j-1}:G_{j}]$ and the last equality, the equation (7) for $M_{\omega}$ becomes:
$M_{\omega}=\sum_{i=q}^{p}k_{i}M_{i}+r_{d}(\ell_{d}-1)M_{\rho_{d}}+\left(\sum_{0}^{q-1}k_{i}m_{i}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\right)\;.$ (8)
If $d=1$, one has $m_{i}=M_{i}$ for $i=0,\ldots,q-1$ and the proof is finished. Otherwise, we apply the induction hypothesis to the last two summands in (8) (in brackets). Notice that the factor $r_{d}\geq 1$ does not affects the proof above and the terms $e_{\rho_{j}}(\varphi_{q})$ (corresponding to a curvette at $E_{\sigma_{q}}$) could be treated (arithmetically) as we did with $e_{\rho_{j}}(\varphi_{\omega})$ for $j<d$. As a consequence, we have:
$M_{\omega}$ $=\sum_{i:\rho_{1}<\tau_{i}\leq\tau_{p}}k_{i}M_{i}+\sum_{i:\tau_{i}\leq\rho_{1}}k_{i}m_{i}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
$=\sum_{i=0}^{{p}}k_{i}M_{i}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
for some integers $r_{j}\geq 1$ ($1\leq j\leq d$). ∎
Remark 5.
For $\sigma\in\check{\Gamma}$, $\tau_{k}\leq\sigma<\tau_{k+1}$ it is known that $m_{\sigma}\in\langle m_{\sigma_{0}},\ldots,m_{\sigma_{k}}\rangle$ or equivalently that $e_{k}$ divides $m_{\sigma}$ (see [12]. Taken into account the expression for $M_{i}$ given in equation (5), this implies that $e_{i-1}$ divides $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{i}}$ and so $\gcd(M_{0},\,\ldots,M_{i})=\gcd(m_{0},\,\ldots,m_{i})=e_{i}$ and
$N_{i}=e_{i-1}/e_{i}=\gcd(M_{0},\,\ldots,M_{i-1})/\gcd(M_{0},\,\ldots,M_{i})\,.$
Lemma 2.
For $i=1,\ldots,g$ one has $(N_{i}-1)M_{\sigma_{i}}\notin\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$, $N_{i}M_{\sigma_{i}}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$ and $N_{i-1}M_{\sigma_{i-1}}<M_{\sigma_{i}}$. In particular, $\{M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ and, moreover, $S$ is the semigroup of values of a complex analytic irreducible curve.
Proof..
Note that, as a consequence of the Remark 5, $(N_{i}-1)M_{i}\notin\langle M_{0},\ldots,M_{i-1}\rangle$.
Let us prove that for $i=1,\ldots,g$, one has that $N_{i-1}M_{i-1}<M_{i}$. If $\rho_{d}\leq\tau_{i-1}<\tau_{i}$, the equation (5) for $M_{i}$ together with the equality $e_{\rho_{i}}=N_{i-1}e_{\rho_{i}}(\varphi_{i-1})$ for $i\leq d$ implie that
$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}=m_{i}+N_{i-1}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}>$
$>N_{i-1}\left(m_{i-1}+\sum_{i=j}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
If $\tau_{i-1}<\rho_{i}\leq\tau_{i}$ the equation (5) for $M_{i-1}$ is $M_{i-1}=m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}$ for some $d^{\prime}<d$. Thus, using the equality
$L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}=\ell_{d-1}L^{d-1}_{j}=\ell_{d-1}\cdots\ell_{d^{\prime}}L^{d^{\prime}}_{j}$
one gets
$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)$
$>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
It is known that the proved properties implies that the semigroup $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ is the semigroup of a complex irreducible plane curve (see [1]).
Proposition 1.
The set of integers $\{M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of the semigroup of values $S^{\mathbb{K}}_{C}$ of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
Proof..
By Lemma 2 it is enough to prove that $S^{\mathbb{K}}_{\nu}$ is generated by $M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}$. It is clear that the semigroup $S^{\mathbb{K}}_{\nu}$ is generated by the integers $M_{\sigma}$, for all $\sigma\in\check{\Gamma}$. Thus one has to show that $M_{\sigma}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ for any $\sigma$.
Let $[\alpha,\beta]$ denote the geodesic in $\check{\Gamma}$ between the vertices $\alpha$ and $\beta$. If $\omega\in[\sigma_{0},\delta_{C}]$, then Lemma 1 gives
$M_{\omega}=\sum_{i=0}^{h}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
where $\rho_{d}<\omega\leq\rho_{d+1}$, $\tau_{h}\leq\omega<\tau_{h+1}$ and $r_{j},k_{i}$ are non-negative integers. The same equation applied to $\rho_{j}$ gives us that $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{i}\rangle$ if $\tau_{i}\leq\rho_{j}<\tau_{i+1}$. In particular, $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{h}\rangle\subset\langle M_{0},\,\ldots,M_{g}\rangle$
If $\omega\in[\sigma_{i},\tau_{i}]$ for some $i=1,\ldots,g$, then $\nu(\varphi_{\omega})=km_{\sigma_{i}}$ for a positive integer $k$. For the conjugate $g\omega$, $g\in G$, one has that $g{\omega}\in[g{\sigma}_{i},g{\tau}_{i}]$ and so $\nu(g{\varphi_{\omega}})=\nu_{g{C}}(\varphi_{\omega})=km_{g{\sigma}_{i}}$. Thus, in this case one has that $M_{\omega}=\nu(\sum g\varphi_{\omega})=k(\sum m_{g\omega})=kM_{\sigma_{i}}$. Note that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$.
Proof..
(Of Theorem 1) It follows directly from Proposition 1 and Lemma 2 taking into account that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$ for all $i\geq 1$. ∎
Remark 6.
As we have proved, $S^{\mathbb{K}}_{\nu}$ is the semigroup of a complex analytic branch $C^{\prime}$ (and so $P^{S}_{C}(t)$ is the usual Poincaré serie of it) different from $C$ if $m_{\rho_{1}}<m_{\tau_{g}}$. We do not know a geometric reason for the existence of the curve $C^{\prime}$.
6 The classical Poincaré series
Theorem 2.
In Cases I and II, one has
$P_{C}(t)$ $=$ $P_{C}^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
$=$ $\frac{\prod_{i=0}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
($s=\infty$ in Case II).
Proof..
Let $P_{C}(t)=\sum_{v=0}^{\infty}a_{v}t^{v}$. For a collection $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$, $k_{\sigma}\in\mathbb{Z}_{\geq 0}$, let ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ be the set of germs $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that the intersection multiplicity of the strict transform of the zero-level set $\{f=0\}$ with the component $E_{\sigma}$ of the exceptional divisor $D$ is equal to $k_{\sigma}$ and, moreover, this strict transform intersects $D$ only at smooth points of the total transform $\pi^{-1}(C)$ of the curve $C$. (Pay attention that the intersection multiplicity of the strict transform of $\{f=0\}$ with the component $E_{g\sigma}$, $g\in G$, is also equal to $k_{\sigma}$.) Let $\nu(\{k_{\sigma}\}):=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$. One can see that $\nu(\{k_{\sigma}\})=\nu(f)$ for any $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.
Remark 7.
Without loss of generality, we shall use the assumption that, for a fixed $v=\nu(\{k_{\sigma}\})$, the latter condition holds automatically. This can be achieved if one makes sufficiently many additional blow-ups at intersection points of the components of the total transform $\pi^{-1}(C)$.
Let $F^{\{k_{\sigma}\}}$ be the image of ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ in the quotient $J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)\cong\mathbb{C}$ with $v=\nu(\{k_{\sigma}\})$. Insted of writing that $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ (and explaining which $\{k_{\sigma}\}$ is considered), we shall sometimes write that $f$ corresponds to the representation $v=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$ or simply to the expression $\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$.
One has:
  1. 1)
    $\bigcup_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}F^{\{k_{\sigma}\}}=J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1);$
  2. 2)
    for each collection $\{k_{\sigma}\}$, $F^{\{k_{\sigma}\}}$ is the complement to a (finite) arrangement of (proper) vector $\mathbb{K}$-subspaces in the vector $\mathbb{K}$-space $\overline{F}^{\{k_{\sigma}\}}$ (the linear span of $F^{\{k_{\sigma}\}}$).
Let $d^{\{k_{\sigma}\}}$ be the dimension (over $\mathbb{K}$) of the vector space $\overline{F}^{\{k_{\sigma}\}}$. One obviously has
$a_{v}=\max_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
For $j=1,2,\ldots,s+1$ ($s$ may be equal to +$\infty$+ ∞), let $A_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic $[\rho_{j},\delta_{C}]$ from $\rho_{j}$ to $\delta_{C}$ (including the ends). For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty. Let the series $P^{(j)}_{C}(t)=\sum_{v=0}^{\infty}a^{(j)}_{v}t^{v}$ be defined by
$a^{(j)}_{v}=\max_{\{k_{\delta}\}\in A_{j}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
One obviously has $P_{C}(t)=P^{(s+1)}_{C}(t)$ if $s<+\infty$ and, if $s=+\infty$ (i. e., if the process of the $\mathbb{K}$-resolution of the curve $C$ is infinite), $P_{C}(t)$ is, in a natural sense, the limit of $P^{(j)}_{C}(t)$ for $j\to+\infty$.
For a function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$, let $\widetilde{f}=f\circ\pi$ be the lifting of $f$ to the surface of the resolution. In order to compute (or evaluate) the dimension $d^{\{k_{\sigma}\}}$, we shall regularly use the following construction. Let $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ be a fixed function with $\nu(h)=\nu(\{k_{\sigma}\})$. (In fact we shall take a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.) For each function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ we shall consider the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$. The function $\Psi$ is regular and different from zero at the intersection point of the exceptional divisor with the strict transform of the curve $C$. The set of values of these ratios for all $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ is isomorphic to the set $F^{\{k_{\sigma}\}}$. This follows from the following arguments. The set $F^{\{k_{\sigma}\}}$ is just the set of the coefficients $a(f)$ in Equation (3) for $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. The coefficient $a(f)$ is equal to the coefficient $a(h)$ (a fixed number) mutiplied by the value of the ratio $\Psi$.
If $\rho_{1}\neq\sigma_{0}$, the statement follows from the following six lemmas (and two Propositions). The necessary changes in the case $\rho_{1}=\sigma_{0}$ will be discussed later.
Lemma 3.
$a^{(1)}_{v}\neq 0$ if and only if $v\in S^{\mathbb{K}}_{C}$.
Proof..
Obviously $a^{(1)}_{v}=0$ for $v\not\in S^{\mathbb{K}}_{C}$. Any $v\in S^{\mathbb{K}}_{C}$ can be represented in the form $v=\sum_{i=0}^{g}k_{i}M_{\sigma_{i}}$ (Proposition 1). Any function $f$ corresponding to this representation gives a non-zero element in the corresponding $F^{\{k_{\sigma}\}}$. Therefore $d^{\{k_{\sigma}\}}\geq 1$ and $a^{(1)}_{v}\geq 1$. ∎
Lemma 4.
For each $v$, $a^{(1)}_{v}\leq 1=[\mathbb{K}_{0}:\mathbb{K}_{0}]$. As a consequence, $a^{(1)}_{v}=1=[\mathbb{K}_{0}:\mathbb{K}_{0}]$ if and only if $v\in S^{\mathbb{K}}_{C}$.
Proof..
Let $\{k_{\sigma}\}\in A_{1}$ and let $h$ be a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$. For $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{1}}$ and $E_{g\delta_{C}}$ for all $g\in G=G_{0}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G$ and therefore belongs to $\mathbb{K}$. ∎
Lemmas 3 and 4 obviously imply the following statement.
Proposition 2.
One has
$P^{(1)}_{C}(t)=P^{S}_{C}(t)=\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\;.$
Lemma 5.
For each $j\geq 1$ and any $v$, $a^{(j)}_{v}\leq[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$.
Proof..
The proof is essentially the same as the one of Lemma 4. For $\{k_{\sigma}\}\in A_{j}$ and a fixed function $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ ($f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$) has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{j}}$ and $E_{g\delta_{C}}$ for all $g\in G_{j-1}$. Therefore this is a (non-zero) constant on them. This constant is invariant with respect to $G_{j-1}$ and therefore belongs to $\mathbb{K}_{j-1}$. ∎
Lemma 6.
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation, assume that $\ell\leq\ell_{i}-1$. We also suppose that $v$ is not of the form $M_{\varkappa}+b^{\prime}$ with $\varkappa>\rho_{j}$ and $b^{\prime}\in S^{\mathbb{K}}_{C}$. (The corresponding case will be anayzed in Lemma 8.) Then $a_{v}^{(j+1)}=\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b}^{(j)}$.
Proof..
We shall use the coordinate $w$ on $E_{\rho_{j}}$ (from the corresponding standard coordinate chart) such that $w=\infty$ at the intersection point with the previous component on the the geodesic between $\sigma_{0}$ and $\rho_{j}$. Let $w=z_{0}\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ be a point on $E_{\rho_{j}}$ blown-up in the process of the resolution of the curve $C$. One has $\mathbb{K}_{j}=\mathbb{K}_{j-1}[z_{0}]$ and $1$, $z_{0}$, $z_{0}^{2}$, …, $z_{0}^{\ell_{j}-1}$ is a basis of the vector space $\mathbb{K}_{j}$ over $\mathbb{K}_{j-1}$.
Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\rho_{j}}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\rho_{j}$, $k^{2}_{\rho_{j}}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$. For certainty, as $h_{1}$ we shall take the $\ell^{\prime}$-th power of the left hand side of the equation of the $G_{j-1}$-curvette at the point $w=1$ of the component $E_{\rho_{j}}$. Let $h=h_{1}h_{2}$. One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesics between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j-1}$ (including both end points of the geodesic). Its value on them belongs to $\mathbb{K}_{j-1}$ and the maximal possible dimension (over $\mathbb{K}$) of the set of its values is $a_{b}^{(j)}$. The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j}$ excluding $\rho_{j}$ itself. Its restriction to the component $E_{\rho_{j}}$ is equal to
$\Psi_{1}|_{E_{\rho_{j}}}=\frac{c\left(\prod\limits_{p=1}^{\ell^{\prime}}(w-w_{p})\right)}{(w-1)^{\ell^{\prime}}}\,,$
where $w_{1},\ldots,w_{\ell^{\prime}}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\rho_{j}}$. All these intersection points $w_{p}$ belong to $\mathbb{K}_{j-1}$. This follows from the fact that the set of intersection points of the strict transform of the curve $\{f_{1}=0\}$ with $E_{\rho_{j}}$ is $G_{j}$-invariant, whereas the orbit of a point not belonging to $\mathbb{K}_{j-1}$ consists of $\ell_{j}$ points. Moreover, its value at $w=\infty$, that is $c$, belongs to $\mathbb{K}$. (This follows from the fact that this ratio is constant on the union of the components from the geodesic $[\sigma_{0},\rho_{j})$.) Thus up to the costant $(z_{0}-1)^{\ell^{\prime}}$ (which does not influence the dimensions) the value of $\Psi$ on the union of the components from the geodesic $(\rho_{j},\delta_{C}]$ is constant and is equal to $c\prod_{p=1}^{\ell^{\prime}}(z_{0}-w_{p})=c(z_{0}^{\ell^{\prime}}-(w_{1}+\ldots+w_{\ell^{\prime}})z_{0}^{\ell^{\prime}-1}+\ldots\pm(w_{1}\cdot\ldots\cdot w_{\ell^{\prime}}))$. Therefore the value of the ratio $\Psi$ on this union is equal to $c(\alpha_{\ell^{\prime}}z_{0}^{\ell^{\prime}}+\alpha_{\ell^{\prime}-1}z_{0}^{\ell^{\prime}-1}+\alpha_{1}z_{0}+\alpha_{0})$, where $\alpha_{0}$, …, $\alpha_{\ell^{\prime}-1}$ are generic points of $\mathbb{K}_{j-1}$ (because we can choose arbitrarily $w_{1},\,\ldots,w_{\ell^{\prime}}$ in $\mathbb{K}_{j-1}$) and $\alpha_{\ell^{\prime}}$ (also a point of $\mathbb{K}_{j-1}$) is the value of the ratio $\Psi_{2}$. The (maximal) dimension of the set of values of $\alpha_{\ell^{\prime}}$ is $a_{b^{*}}^{(j)}$, where $b^{*}=v-\ell^{\prime}M_{\rho_{j}}$. Since $1$, $z_{0}$, …$z_{0}^{\ell^{\prime}}$ are linear independent over $\mathbb{K}_{j-1}$, the total dimension is equal to $\ell^{\prime}[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b^{*}}^{(j)}$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a^{(j)}_{b}$.
Lemma 7.
If $\rho_{j}<\varkappa\leq\delta_{C}$ (that is $\varkappa$ is on the geodesic between $\rho_{j}$ and $\delta_{C}$ and $\varkappa\neq\rho_{j}$), then $a_{M_{\varkappa}}^{(j+1)}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$.
Proof..
Any collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=M_{\varkappa}$ obviously belongs to $A_{j+1}$. According to Lemma 5, $a_{M_{\varkappa}}^{(j+1)}\leq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. Let us take the collection $\{k^{0}_{\sigma}\}$ with $k^{0}_{\varkappa}=1$ and $k^{0}_{\sigma}=0$ for $\sigma\neq\varkappa$. One has $\nu(\{k^{0}_{\sigma}\})=M_{\varkappa}$. Let $f=0$ be an equation of a $G_{\varkappa}(=G_{j})$-curvette intersecting the component $E_{\varkappa}$ at the point with the coordinate $\alpha\in\mathbb{K}_{j}$. Here we assume that the coordinate is equal to zero at the intersection point with the component on the geodesic to $\delta_{C}$ and to $\infty$ at the one with the component on the geodesic to $\sigma_{0}$. Let $g$ be one of these functions, say, with $\alpha=1$. The ratio $\Psi={\widetilde{f}}/{\widetilde{g}}$, for $f\in{\mathcal{E}}^{\{k^{0}_{\sigma}\}}_{\mathbb{K}^{2},0}$, has no zeroes or poles on all the components of each of the indicated geodesics and thus is constant on each of them. On the component $E_{\varkappa}$ this ratio is of the form $c\frac{w-\alpha}{w-1}$ with $c\in\mathbb{K}$ and (arbitrary!) $\alpha\in\mathbb{K}_{j}$. The value on $\delta_{C}$ (equal to the value at $w=0$, i.e. $c\alpha$) is an arbitrary point of $\mathbb{K}_{j}$. Therefore $d^{\{k^{0}_{\sigma}\}}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$ and $a_{M_{\varkappa}}^{(j+1)}\geq[\mathbb{K}_{j}:\mathbb{K}_{0}]$. ∎
Lemma 8.
Let $\varkappa$ be as in Lemma 7. Due to Lemma 1, one has that $M_{\varkappa}=(\ell_{j}-1)M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$. Then $a_{b}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$.
Proof..
Due to Lemma 1, one has $M_{\varkappa}=b^{\prime}+\sum_{q=1}^{j}(\ell_{q}-1)M_{\rho_{q}}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$. Let $c_{k}:=b^{\prime}+\sum_{q=1}^{k}(\ell_{q}-1)M_{\rho_{q}}$, $k=1,\ldots,j$. We have to show that $a^{(j)}_{c_{j-1}}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$. We shall show that $a^{(k)}_{c_{k-1}}=[\mathbb{K}_{k-1}:\mathbb{K}_{0}]$ for all $k$. We shall use induction in $k$. For $k=1$ this is trivial: $a^{(1)}_{c_{0}}=1=[\mathbb{K}_{0}:\mathbb{K}_{0}]$. Due to Lemma 6, we have $a^{(k)}_{c_{k-1}}=(\ell_{k-1}-1)[\mathbb{K}_{k-2}:\mathbb{K}_{0}]+a^{(k-1)}_{c_{k-2}}$. By the induction assumption $a^{(k-1)}_{c_{k-2}}=[\mathbb{K}_{k-2}:\mathbb{K}_{0}]$ and therefore $a^{(k)}_{c_{k-1}}=\ell_{k-1}[\mathbb{K}_{k-2}:\mathbb{K}_{0}]=[\mathbb{K}_{k-1}:\mathbb{K}_{0}]$. ∎
Lemmas 58 imply the following statement.
Proposition 3.
One has
$P^{(j+1)}(t)$ $=$ $P^{(j)}(t)\cdot(1+t^{M_{\rho_{j}}}+t^{2M_{\rho_{j}}}+\ldots+t^{(\ell_{j}-1)M_{\rho_{j}}})$
$=$ $P^{(j)}(t)\cdot\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}.$
Proof..
Equation (3) is equivalent to
$a_{v}^{(j+1)}=a_{v}^{(j)}+a_{v-M_{\rho_{j}}}^{(j)}+a_{v-2M_{\rho_{j}}}^{(j)}+\ldots+a_{v-(\ell_{j}-1)M_{\rho_{j}}}^{(j)}\,.$ (10)
If $v\in S^{\mathbb{K}}_{C}$ cannot be represented as $M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, one has $a_{v}^{j+1}=a_{v}^{j}$, what is compatible with (10). Let $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, $11\leq\ell<\ell_{j}-$ 1, and $v$ cannot be represented as $(\ell+1)M_{\rho_{j}}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$. In this case Equation (10) reduces to
$a_{v}^{(j+1)}=a_{v}^{(j)}+a_{v-M_{\rho_{j}}}^{(j)}+a_{v-2M_{\rho_{j}}}^{(j)}+\ldots+a_{v-(\ell-1)M_{\rho_{j}}}^{(j)}+a_{b}^{(j)}\,$ (11)
(all other summands are equal to zero). Since, for $10\leq\ell^{\prime}\leq\ell-$ 1, $v-\ell^{\prime}M_{\rho_{j}}=M_{\rho_{j}}+b^{*}$ with $b^{*}\in S^{\mathbb{K}}_{C}$ and, due to Lemma 7, $a_{M_{\rho_{j}}}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$, all the summands in (11) except the last one are equal to $[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$. Thus in this case (10) follows from Lemma 6.
Let $v=(\ell_{j}-1)M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$. There are two possibilities.
  1. 1)
    The value $v$ can be represented as $M_{\varkappa}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$, $\varkappa>\rho_{j}$. In this case (10) follows from Lemmas 7 and  8.
  2. 2)
    The value $v$ cannot be represented in the indicated form. In this case (10) follows from Lemma 6 again.
Remark 8.
One can see that the proof of Proposition 3 gives a little bit more strong statement which we shall use in Section 7. Let $A^{\prime}_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic from $\rho_{j}$ to $\delta_{C}$ excluding (!) $\rho_{j}$ and including $\delta_{C}$. For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty. Let the series $\widetilde{P}^{(j)}_{C}(t)=\sum_{v=0}^{\infty}\widetilde{a}^{(j)}_{v}t^{v}$ be defined by
$\widetilde{a}^{(j)}_{v}=\max_{\{k_{\delta}\}\in A^{\prime}_{j}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
Then one has
$\widetilde{P}^{(j)}_{C}(t)=P^{(j+1)}(t)=P^{(j)}(t)\cdot\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}.$ (12)
Propositions 2 and 3 imply the statement of Theorem 2 in the case under consideration: $\rho_{1}\neq\sigma_{0}$. In the case $\rho_{1}=\sigma_{0}$, Lemmas 58 and thus Proposition 3 hold for $j\geq 2$. Proposition 2 does not hold. However, it is possible to compute the series $P^{(2)}_{C}(t)$ directly using Lemmas 58. Thus the statement of Theorem 2 holds in this case as well. ∎
Remark 9.
Any series $\zeta(t)=1+c_{1}t+c_{2}t^{2}+\cdots\in\mathbb{Z}[[t]]$ can be in a unique way represented as $\zeta(t)=\prod_{m=1}^{\infty}(1-t^{m})^{s_{m}}$ with $s_{m}\in\mathbb{Z}$. We say that the series is cyclotomic if this product/ratio of binomials is finite. There is a natural generalization of this notion for series in sevearl variables. Ussually in previous computations (say e.g. [5]) the Poincaré series appeared to be cyclotomic. This is not the case for the classical Poincaré series $P_{C}(t)$ in case II, i.e. when $s=\infty$.
Remark 10.
Let us assume that the curve valuation is of type I. Let $c$ be the conductor of the semigroup $S^{\mathbb{K}}_{C}$ (i.e the minimal element of $S^{\mathbb{K}}_{C}$ such that $c+\mathbb{Z}_{\geq 0}\subset S^{\mathbb{K}}_{C}$). It is known that $c$ can be expresed as
$c=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1\;.$
Let $\Delta:=c+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}$ and $\ell=\ell_{1}\cdots\ell_{s}=[\mathbb{K}_{s}:\mathbb{K}]$. The computation of the Poincaré series gives that $a_{v}=\dim J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)=\ell$ for all $v\geq\Delta$ and $a_{\Delta-1}<\ell$. Moreover, one has the following symmetry property for all $v\in\mathbb{Z}$:
$\dim\left({J^{\mathbb{K}}(v)}\left/{J^{\mathbb{K}}(v+1)}\right.\right)+\dim\left({J^{\mathbb{K}}(\Delta-1-v)}\left/{J^{\mathbb{K}}(\Delta-v)}\right.\right)=\ell\;$ (13)
Following the notations of Remark 2, the integral closure $\overline{R}$ of the local ring $R=\mathbb{K}[[x,y]]/(F)$ of the algebroid curve defined by $F$ is a valuation ring and is isomophic to $\mathbb{K}^{\prime}[[u]]$ ($u$ is an uniformizing parameter). One has:
Corollary 1.
Let
$\Delta=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}\;.$
Then, the conductor ideal ${\mathcal{C}}(\overline{R}/R)=\{z\in\overline{R}:z\overline{R}\subset R\}$ is the ideal $u^{\Delta}\overline{R}$.
Note that the ring $R$ is Gorenstein (being plane) and the symmetry of the equation (13) is the just the characterization of the Gorenstein condition of the ring $R$ (cf. [3]).
7 The Poincaré series of a divisorial valuation
Let $\nu=\nu_{\delta}$ be a divisorial valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$. This means that it is defined by a component $E_{\delta}$ of the exceptional divisor of a modification of $(\mathbb{C}^{2},o)$. Let $\pi:(W,{\mathcal{D}})\to(\mathbb{C}^{2},0)$ be the minimal $\mathbb{K}$-resolution of the valuation $\nu$. The quotient of the dual graph of this resolution by the action of the Galois group $G$ looks like in figure 6.
Figure 6: The resolution graph $\check{\Gamma}$ for a divisorial valuation.
This graph essentially coincides with the minimal resolution graph of a $\mathbb{K}_{\delta}$-curvette at the component $E_{\delta}$ ($\mathbb{K}_{\delta}=\mathbb{K}_{s}$) with a possible tail from $\tau_{g}$ to $\delta$ added. Here, as above, $\sigma_{i}$, $i=0,1,\ldots,g$ are the dead ends of the graph $\check{\Gamma}$, $\tau_{i}$, $i=1,\ldots,g$ are the rupture points and $\rho_{j}$, $j=1,\ldots,s$ are the branching points of the $\mathbb{K}$-resolution.
Remark 11.
Pay attention that $\rho_{s}\neq\delta$. One may define $\rho_{s+1}$ as $\delta$. In this case the divisorial valuation $\nu_{\delta}$ coincides with the curve valuation $\nu_{C^{*}}$ for $C^{*}$ being a curvette at the component $E_{\delta}$ at a non-algebraic point of it. (Thus $\ell_{s+1}$ is assumed to be equal to $\infty$.)
We shall keep the notations ${\mathcal{E}}^{\{k_{\sigma}\}}_{\mathbb{K},0}$, $F^{\{k_{\sigma}\}}$, and $d^{\{k_{\sigma}\}}$ as before.
Theorem 3.
One has
$P_{\nu}(t)=P^{S}(t)\frac{1}{1-t^{M_{\delta}}}\cdot\prod_{i=1}^{s}\frac{1-t^{\ell_{i}M_{\rho_{i}}}}{1-t^{M_{\rho_{i}}}}\;.$ (14)
Proof..
Let $C$ be a $\mathbb{K}_{s}$-curvette at the component $E_{\delta}$. For the coefficient $a_{v}$ in the Poincaré series $P_{\nu}(t)=\sum a_{v}t^{v}$ one has $a_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}$ such that $\nu(\{k_{\sigma}\})=\sum k_{\sigma}M_{\sigma}=v$. Let $A^{\prime}$ be the set of collections $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$ such that $k_{\delta}=0$ and let $P^{\prime}(t)=\sum a^{\prime}_{v}t^{v}$ with $a^{\prime}_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}\in A^{\prime}$ such that $\nu(\{k_{\sigma}\})=v$.
Lemma 9.
One has
$P^{\prime}_{\nu}(t)=P_{C}(t)=P^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}\;.$
Proof..
The statement is a direct consequence of Remark 8. ∎
Lemma 10.
Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\delta}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation. Then $a_{v}=\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$.
Proof..
(Cf. proof of Lemma 6.) Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\delta}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\delta$, $k^{2}_{\delta}=0$. A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$. Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $h=h_{1}h_{2}$. As in Lemma 6, we shall use a coordinate on $E_{\delta}$ such that the intersection point with the previous component from the geodesic to $\sigma_{0}$ is equal to $\infty$ and we shall take the function $h_{1}$ described there (that is, the $\ell^{\prime}$th power of the left hand side of the equation of the $G_{s}$-curvette at the point $w=1$ of the component $E_{\delta}$). One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$. The function $\Psi_{2}$ is constant on the component $E_{\delta}$. The maximal possible dimension (over $\mathbb{K}$) of the set of its values on this component is $a^{\prime}_{b^{\prime}}$ ($b^{\prime}=v-\ell^{\prime}M_{\delta}$). The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\sigma_{0}$ and $\delta$, excluding the last one. Its restriction to the component $E_{\delta}$ is equal to $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})/(w-1)^{\ell^{\prime}}$ with $c\in\mathbb{K}$, where $w_{p}\in\mathbb{K}_{s}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\delta}$. Therefore, up to the fixed multiple $(w-1)^{\ell^{\prime}}$, the restriction of $\Psi$ to the component $E_{\delta}$ is equal to the restriction of $\Psi_{2}$ multiplied by $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})$. The dimension (over $\mathbb{K}$) of the space of functions of the latter form is $\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}]$ (since $w_{p}$ are arbitrary elements of $\mathbb{K}_{s}$). Therefore the (maximal) dimension of the set of these functions on $E_{\delta}$ is $a^{\prime}_{b^{\prime}}+\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}_{0}]$. Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$. ∎
Just in the same way as Proposition 3 we get the following statement.
Proposition 4.
One has
$P_{\nu}(t)=P^{\prime}_{\nu}(t)\cdot\frac{1}{1-t^{M_{\delta}}}.$
Lemma 9 and Proposition 4 imply the statement of Theorem 3. ∎
Remark 12.
One can see that the classical Poincaré series of a curve or of a divisorial valuation determines the combinatorial type of the minimal $\mathbb{K}$-resolution of the valuation.

### v2

@@
 Abstract
-For a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers, we consider curve and divisorial valuations on the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables with the coeficients in $\mathbb{K}$.
-We compute the semigroup Poincaré series and the classical Poincaré series of a valuation.
+For a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers, we consider curve and divisorial valuations on the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables with the coefficients in $\mathbb{K}$.
+We compute the semigroup Poincaré series and the classical Poincaré series of a (discrete, rank one) valuation on it.
 1 Introduction
@@
 The Poincaré series of one divisorial valuation was computed in [12].
-The Poincaré series of a collection of curve valuations was computed in [5].
-It was shown that this series coincides with the Alexander polynomial in several variables of the corresponding algebraic link.
+The Poincaré series of a collection of curve valuations was computed in [4] (see also [5] for a “simpler” proof).
+It was shown that (for a collection consisting of more than one valuation) this series coincides with the Alexander polynomial in several variables of the corresponding algebraic link.
 This implies that the Poincaré series determines the topology of the union of the corresponding curves.
@@
 In [8], there was computed the Poincaré series of one divisorial valuation in this setting.
+Here we consider Poincaré series of valuations on the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$ of germs of holomorphic functions in two variables whose Taylor coefficients are from a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers (or, equivalently, on the algebra $\mathbb{K}[[x,y]]$ of formal power series).
+We compute the Poincaré series of one curve or one divisorial valuation on it: Theorems 1, 2, and 3.
+The Poincaré series are expressed in terms of a resolution of a special form and are represented as products/ratios of binomials of the form $(1-t^{m})$.
+The computation of the Poincaré series of a curve valuation presented in [13] was combinatorial.
+However, it was known long before that.
+The importance of that paper was the observation that it coincided with the monodromy zeta function, i.
+e., with the Alexander polynomial of the corresponding algebraic knot divided by (1t)(1-t).
+All the computations of the Poincaré series made by the authors (starting from [4]) used some elements of topology: versions of the integration with respect to the Euler characteristic.
+The key property for that was the fact that the (additive) Euler characteristic of a complex affine space is equal to $1$.
+This does not work already for real analogues of the Poincaré series in [7] and [8] (the Euler characteristic of a real affine space is equal to $\pm 1$) and makes no (straightfoward) sense for affine spaces over subfields of $\mathbb{C}$.
+Therefore the computations in these cases (at least at the contemporary stage) are essentially combinatorial (modulo a certain analytic hint explained in Section 6).
+The content of the paper is the following.
+In Section 2, we discuss curve valuations on the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
+Section 3 is devoted to a description of a resolution process of an irreducible curve playing the central role in the computations.
+In Section 4, one associates Galois groups and subfields to the components of the exceptional divisor of a resolution.
+Section 5 is devoted to the computation of the semigroup Poincaré series, i.
+e., the sum $\sum_{v\in S^{\mathbb{K}}_{\nu}}t^{v}$ over the set $S^{\mathbb{K}}_{\nu}$ of values of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$: the semigroup of values of the valuation.
+In Section 6, the classical Poincaré series (1) of a curve valuation is computed.
+In Section 7 it is computed for a divisorial valuation.
+We are very thankful to the referee for reading carefully the paper and making a number of useful suggestions.
 2 Curve valuations on functions over a field
@@
 We shall refer to ${\mathcal{E}}_{\mathbb{K}^{2},0}$ because of some convenience of the language.
-Let $(C,0)$ be a plane algebroid branch on $(\mathbb{C}^{2},0)$, that is $C$ (after an appropriate change of the coordinates in $\mathbb{C}^{2}$) is defined by
-$x=x(\tau)=\tau^{m},\\ y=y(\tau)=\sum_{i\geq m}c_{i}\tau^{i}\in\mathbb{C}[[\tau]]\,.$ (2)
-(We assume that the parametrization (2) is irreducible, i.
+Remark 1.
+Strictly speaking, some constructions below (in particular the description of the action of the Galois group $G$ on branches at the beginning of section 3) cannot be applied directly to the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
+The series for the branch $g\gamma$ and for the function $gh$ therein could be non-convergent.
+These constructions work in the algebra $K[[x,y]]$, however use of it everywhere produces some language problems.
+We do not pay attention to this problem since, for a bounded values of valuations we can treat not functions or series, but jets of order high enough, what means essentially polynomials.
+Informally one can say that the computations can be made in a sort of “an algebra of jets”.
+Let $C$ be a plane algebroid branch on $(\mathbb{C}^{2},0)$, that is $C$ (after an appropriate change of the coordinates in $\mathbb{C}^{2}$ defined over $\mathbb{K}$) is given by
+$x=x(\tau)=\tau^{m},\\ y=y(\tau)=\sum_{i\geq m}c_{i}\tau^{i}\in\mathbb{C}[[\tau]]\,$ (2)
+with some $m>0$.
+(We assume that the parametrization (2) is primitive, i.
 e., the greatest common divisor of $m$ and of all $i$ such that $c_{i}\neq 0$ is equal to $1$.) The branch $C$ defines a valuation $\nu=\nu_{C}$ on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of holomorphic function germs in $x$ and $y$.
@@
 Definition 1.
-The $\mathbb{K}$-Poincaré series of the branch $C$ is
-$P_{C}(t)(=P_{C}^{\mathbb{K}}(t))=\sum_{v=0}^{\infty}\dim_{\mathbb{K}}\left(J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\right)\cdot t^{v}.$ (4)
-Example 1.
-If $\mathbb{K}=\mathbb{R}$, $P_{C}(t)$ is the classical Poincaré series of the valuation $\nu$ computed in [7].
+The (classical) $\mathbb{K}$-Poincaré series of the branch $C$ is
+$P_{C}^{\mathbb{K}}(t)=\sum_{v=0}^{\infty}\dim_{\mathbb{K}}\left(J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\right)\cdot t^{v}.$ (4)
+If $\mathbb{K}=\mathbb{R}$, $P^{\mathbb{R}}_{C}(t)$ is the classical Poincaré series of the valuation $\nu$ computed in [7].
 Remark 2.
-If $\mathbb{K}=\mathbb{C}$, all the coefficients of the series (4) are equal to 00 or $1$.
+If $\mathbb{K}=\mathbb{C}$, all the coefficients of the series (4) are equal to 0 or 11.
 This is not the case, in general, for a proper subfield $\mathbb{K}\subset\mathbb{C}$.
-In particular, if $\mathbb{K}=\mathbb{R}$ (the field of real numbers), all the coefficients are equal to 00, $1$, or $2$ ([7]).
-Let $S^{\mathbb{K}}_{C}:=\{\nu(f):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}\}\subset\mathbb{Z}_{\geq 0}$ be the semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
+In particular, if $\mathbb{K}=\mathbb{R}$ (the field of real numbers), all the coefficients are equal to 0, 11, or 22 ([7]).
+Let $S^{\mathbb{K}}_{C}:=\{\nu_{C}(f):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}\}\subset\mathbb{Z}_{\geq 0}$ be the semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
 Definition 2.
 The $\mathbb{K}$-semigroup Poincaré series of the valuation $\nu$ is $P^{S}_{C}(t)=\sum_{v\in S^{\mathbb{K}}_{C}}t^{v}$.
+For $\mathbb{K}=\mathbb{R}$ this is the semigroup Poincaré series considered (and computed) in [7].
 Example 1.
-For $\mathbb{K}=\mathbb{R}$ this is the semigroup Poincaré series considered (and computed) in [7].
-Here we compute the $\mathbb{K}$-semigroup Poincaré series and the $\mathbb{K}$-Poincaré series for one curve or divisorial valuation.
+(cf.
+[7, Example 1 and Remark 3]) Let $\mathbb{K}=\mathbb{R}$, and let the branch $C$ be given by the parametrization $(\tau^{4},\alpha\tau^{6}+\tau^{7})$ with a generic complex $\alpha$.
+The semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ is equal to $S_{C}^{\mathbb{R}}=\langle 4,6,25\rangle$.
+Here $\langle\bullet,\bullet,\ldots\rangle$ denotes the subsemigroup of $\mathbb{Z}_{\geq 0}$ generated by the corresponding natural numbers:
+$\langle m_{1},\ldots,m_{p}\rangle=\{m\in\mathbb{Z}_{\geq 0}:m=k_{1}m_{1}+\ldots+k_{p}m_{p}\text{\ \ with\ \ }k_{i}\in\mathbb{Z}_{\geq 0}\}.$
+This semigroup coincides with the semigroup of values of the curve valuation $\nu_{C^{\prime}}$ on ${\mathcal{O}}_{\mathbb{C}^{2},0}$, where the curve $C^{\prime}$ is given by the parametrization $(\tau^{4},\tau^{6}+\tau^{19})$.
+A direct relation between the curves $C$ and $C^{\prime}$ is not clear.
+The semigroup Poincaré series (both of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ and of the valuation $\nu_{C^{\prime}}$ on ${\mathcal{O}}_{\mathbb{C}^{2},0}$) is given by
+$P^{S}_{C}(t)=\frac{(1-t^{12})(1-t^{50})}{(1-t^{4})(1-t^{6})(1-t^{25})}\,.$
+The classical Poincaré series of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ is given by
+$P^{\mathbb{R}}_{C}(t)=\frac{(1-t^{24})(1-t^{50})}{(1-t^{4})(1-t^{6})(1-t^{25})}\,.$
 3 The resolution process of a branch
-Let $G$ be the Galois group of the extention $\mathbb{C}/\mathbb{K}$.
+Let $G$ be the Galois group of the extension $\mathbb{C}/\mathbb{K}$.
 i.
-e.
-the group of automorphisms of $\mathbb{C}$ which are trivial on $\mathbb{K}$.
-Let us consider the minimal $G$-invariant resolution process of the branch $C$ (by a sequence of blow-ups).
+e., the group of automorphisms of the field $\mathbb{C}$ which are trivial on $\mathbb{K}$.
+The group $G$ acts on branches (irreducible complex plane curve germs).
+This action can be defined by any of the following three ways.
+  1. 1)
+    Let a branch $\gamma$ be given by a parametrization $x=\tau^{k}$, $y=\sum\limits_{i\geq k}a_{i}\tau^{i}$.
+Then the branch $g\gamma$, $g\in G$, is given by the parametrization $x=\tau^{k}$, $y=\sum\limits_{i\geq k}g(a_{i})\tau^{i}$.
+  2. 2)
+    Let a branch $\gamma$ be given by an equation $h(x,y)=0$, where $h(x,y)=\sum\limits_{i,j\geq 0}b_{ij}x^{i}y^{j}$.
+Then the branch $g\gamma$ is given by the equation $gh(x,y)=0$, where $gh(x,y):=\sum\limits_{i,j\geq 0}g(b_{ij})x^{i}y^{j}$.
+  3. 3)
+    Let a branch $\gamma$ be considered as a subset of $(\mathbb{C}^{2},0)$: $(\gamma,0)\subset(\mathbb{C}^{2},0)$.
+The group $G$ acts on the plane $\mathbb{C}^{2}$ by $g(x,y)=(g(x),g(y))$.
+Then the branch $g\gamma$ as a subset of $(\mathbb{C}^{2},0)$ is just the image of $\gamma$ under this action.
+Remark 3.
+For fixed $x$ and $y$ the parametrization of $\gamma$ is not unique.
+One can obtain another parametrization from one of them multiplying the parameter $\tau$ by a root $\xi$ of degree $k$ of $1$.
+For $g\in G$, $g(\xi)$ is also a root of degree $k$ of $1$.
+Thus this only leads to another parametrization the branch $g\gamma$.
+It is not difficult to see that, for a $G$-invariant curve germ $\mathcal{C}$} (possibly reducible), there exists a function germ $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that ${\mathcal{C}}=\{h=0\}$.
+Let us consider the minimal $G$-invariant resolution process of the branch $C$ defining the valuation (by a sequence of blow-ups).
 This means the following.
@@
   1. 1)
-    There are finitely many different branches in the $G$-orbit of the strict transform $\widetilde{C}$ of the branch $C$, all of them are smooth and they intersect the exceptional divisor transversally at pairwise different points.
-We get (the minimal embedded $G$-invariant) resolution of $C$ and the process is finished.
+    There are finitely many different branches in the $G$-orbit of the branch $C$, their strict transforms are smooth and intersect the exceptional divisor transversally at pairwise different points.
+We get the (minimal embedded) $G$-invariant resolution of $C$ and the process is completed.
   2. 2)
-    The branches from the $G$-orbit of $\widetilde{C}$ (whose number may be infinite) intersect the exceptional divisor at finite number of points, but either they are not smooth, or do not intersect the exceptional divisor transversally, or (some) different branches from the orbit intersect it at the same point.
-In this case one has to continue the process.
+    The strict transforms of the branches from the $G$-orbit of $C$ (whose number may be infinite) intersect the exceptional divisor at finite number of points, but either they are not smooth, or do not intersect the exceptional divisor transversally, or some (different) of them intersect it at the same point.
+(In this case, if the $G$-orbit of $C$ has infinitely many branches, the number of strict transforms intersecting the exceptional divisor at one point is also infinite.) In this case one has to continue the resolution process.
   3. 3)
-    The branches from the $G$-orbit of $\widetilde{C}$ intersect the exceptional divisor at infinitely many different points. In this case the process is terminated.
-In Case 2 one has to blow-up (all) the intersection points of the branches from the $G$-orbit of $\widetilde{C}$ with the exceptional divisor (their number being finite).
-One gets one of the analogues the cases 1–3 again.
+    The strict transforms of the branches from the $G$-orbit of C{C} intersect the exceptional divisor at infinitely many different points.
+In this case the process is interrupted.
+(Do not mix with the term “completed”.
+In this case, one cannot get a $G$-invariant resolution since blowing-ups infinitely many points does not produce a complex surface.)
+In Case 2) one has to continue the resolution process blowing-up (all) the intersection points of the strict transforms of the branches from the $G$-orbit of $C$ with the exceptional divisor (their number being finite).
+One gets one of the analogues of the cases 1)–3) again.
 Etc.
-As the result we arrive to one of the following three situations.
-  1. I
-    There are finitely many branches in the $G$-orbit of $C$ and they are resolved at a certain step. The dual graph of the resolution looks like in Figure 1.
+As the result, we arrive to one of the following three situations.
+  1. I.
+    There are finitely many branches in the $G$-orbit of $C$ and they are resolved at a certain step.
+  2. II.
+    The $G$-orbit of $C$ has infinitely many branches, but the described process goes on without interruption.
+This means that at each step one has to blow-up finitely many points, but there are infinitely many steps like that.
+One does not arrive to a $G$-invariant resolution of $C$, but has got a $G$-invariant resolution process (an infinite one).
+  3. III.
+    The process is interrupted at a certain step.
+This means that at this step the strict transforms of the branches from the $G$-orbit of $C$ intersect each of the new-born components of the exceptional divisor at infinitely many points.
+In this case one gets neither a $G$-invariant resolution of $C$, nor a resolution process.
+The cases I–III take place in the following situations.
+  1. I.
+    All the coefficients in the series (2) are from a finite extension of $\mathbb{K}$. This is equivalent to the condition that the $G$-orbit of $C$ has finitely many elements.
+  2. II.
+    The coefficients all together are not from a finite extension of $\mathbb{K}$, but each of them is algebraic over $\mathbb{K}$.
+  3. III.
+    There exists (at least) one coefficient transcendental over $\mathbb{K}$.
+At each step of a $G$-invariant resolution of $C$, the action of the Galois group $G$ on $\mathbb{C}^{2}$ lifts to its action on the surface of modification and thus induces an action on the set of irreducible components of the exceptional divisor.
+These components are in bijection with the vertices of the dual graph of the modification and thus one has an action of $G$ on this dual graph.
+In the case I, the dual graph $\Gamma$ of the resolution looks like in Figure 1.
 Figure 1: The resolution graph $\Gamma$ in case I (here $s=2$, $g=6$, see below).
-The Galois group $G$ acts on the graph $\Gamma$ and the quotient $\check{\Gamma}$ by the action looks like in Figure 2 (and further on).
+For a description of the resolution (or of the resolution process), it is convenient to use the quotient $\check{\Gamma}$ of the graph $\Gamma$ by the action of the Galois group $G$.
+In this case it looks like in Figure 2.
 The graph $\Gamma$ contains a subgraph which geometrically coincides with the dual graph of a usual resolution of the curve $C$: the lower part of $\Gamma$ on Figure 1.
-(We do not consider this resolution below.) We use “short” notations (by greek letters) for verticies on this subgraph (and also for the corresponding verticies of the graph $\check{\Gamma}$).
-All other verticies are obtain from them by the action of elements of the Galois group $G$.
-In Figure 1, $\sigma_{i}$, $i=0,\ldots,g$, (and also their images under the action of $G$) are the dead ends of the graph ($g$ is the number of the Puiseux pairs of the curve $C$), $\tau_{i}$, $i=1,\ldots,g$, (and also their images) are the rupture points of it, $\rho_{j}$, $j=1,\ldots,s$, ($s$ is equal to $\infty$ in Case II) are the splitting points corresponding to the changes of the isotropy subgroups for the $G$-action on $\Gamma$ (that is the splitting points for usual resolution of different branches from the orbit of $C$, there are finitely many of them in Case I infinitely many in Case II.
-Figure 2 shows $\check{\Gamma}$ in the case when the process of the $\mathbb{K}$-resolution of $C$ is not finish when $C$ itself is resolved. Otherwise $\delta_{C}$ coincides with $\tau_{g}$.
-Figure 2: The resolution graph $\check{\Gamma}$ in Case I (with arbitrary $g$ and $s$).
-  1. II
-    The described process goes on without termination.
-This means that we get an infinite sequence of blow-ups.
-The corresponding dual graph of the process looks like in Figure 3.
-One can say that $s$ is equal to infinity in this case.
-    Figure 3: The resolution graph $\check{\Gamma}$ in Case II (with $s=\infty$).
-  2. III
-    The process is terminated at a certain step.
-This means that at this step the branches from the $G$-orbit of $\widetilde{C}$ intersect each of the new-born components of the exceptional divisor at infinitely many points.
-The dual graph of the process looks like in Figure 4.
-    In the latter case, the valuation coincides with $n\nu$, where $\nu=\nu_{\rho_{s}}$ is the divisorial valuation defined by $E_{\rho_{s}}$ and $n$ is the intersection multiplicity $[E_{\rho_{s}},\widetilde{C}]$ of the strict transform $\widetilde{C}$ of the curve $C$ with the component $E_{\rho_{s}}$.
-    Figure 4: The resolution graph $\check{\Gamma}$ in Case III.
-Cases I–III take place in the following situations.
-  1. I
-    All the coefficients in the series (2) are from a finite extension of $\mathbb{K}$. This is equivalent to the condition that the orbit $GC$ has finitely many elements.
-  2. II
-    The coefficients all together are not from a finite extension of $\mathbb{K}$, but each of them is algebraic over $\mathbb{K}$.
-  3. III
-    There exists (at least) one coefficient transcendental over $\mathbb{K}$.
+(We do not consider this resolution below.) We use “short” notations (by greek letters) for vertices on this subgraph (and also for the corresponding vertices of the graph $\check{\Gamma}$).
+All other vertices are obtained from them by the action of elements of the Galois group $G$.
+In Figures 1 and 2, $\sigma_{i}$, $i=0,1,\ldots,g$, (and also their images under the action of $G$) are the dead ends of the graph ($g$ is the number of the Puiseux pairs of the curve $C$), $\tau_{i}$, $i=1,\ldots,g$, (and also their images) are the rupture points of it, $\rho_{j}$, $j=1,\ldots,s$, are the splitting points corresponding to the changes of the isotropy subgroups for the $G$-action on $\Gamma$ (that is the splitting points for usual resolutions of different branches from the orbit of $C$, there are finitely many of them in Case I).
+Let $\delta_{C}$ be the vertex corresponding to the component of the exceptional divisor intersecting the strict transform of the curve $C$.
+In Figure 1, $\delta_{C}=\tau_{6}$.
+Figure 2 shows $\check{\Gamma}$ in the case when the process of the $G$-resolution of $C$ is not finished when $C$ itself is resolved, $\delta_{C}=\rho_{s}$ there.
+(Otherwise $\delta_{C}$ coincides with $\tau_{g}$.)
+Figure 2: The quotient $\check{\Gamma}$ of the resolution graph in Case I (with arbitrary $g$ and $s$).
+In Case II, one has the dual graph of the resolution process. The corresponding quotient $\check{\Gamma}$ looks like in Figure 3. One can say that $s$ is equal to infinity in this case.
+Figure 3: The quotient $\check{\Gamma}$ of the graph of the resolution process in Case II (with $s=\infty$).
+In Case III, the quotient of the modification graph looks like in Figure 4. The curve $C$ may be not resolved at the interruption step.
+In the latter case, the valuation coincides with $n\nu$, where $\nu=\nu_{\rho_{s}}$ is the divisorial valuation defined by $E_{\rho_{s}}$ and $n$ is the intersection multiplicity $E_{\rho_{s}}\circ\widetilde{C}$ of the strict transform $\widetilde{C}$ of the curve $C$ with the component $E_{\rho_{s}}$.
+Figure 4: The quotient $\check{\Gamma}$ of the modification graph in Case III.
+Example 2.
+Let $\mathbb{K}=\mathbb{Q}$ and the branch $C$ given by the parametrization $x=\tau^{4}$, $y=\alpha_{0}\tau^{6}+t^{7}+\sum_{i\geq 1}\alpha_{i}\tau^{7+i}$ where $\alpha_{0},\alpha_{1},\ldots\in\mathbb{C}\setminus\{0\}$.
+Let $\mathbb{K}^{\prime}:=\mathbb{Q}(\alpha_{0},\alpha_{1},\ldots)\subset\mathbb{C}$.
+Figure 5: The dual graph of $C$
+One can distinguish the following cases:
+  1. 1.
+    $\mathbb{K}^{\prime}$ is finite algebraic extension of $\mathbb{Q}$.
+In this case $\mathbb{K}^{\prime}=\mathbb{Q}(\alpha_{0},\,\ldots,\alpha_{k})$ for some $k$ and if we add the conditions $\alpha_{0}\notin\mathbb{Q}$ (algebraic, generic) and $\alpha_{i}\notin\mathbb{Q}(\alpha_{0},\,\ldots,\alpha_{i-1})$ (algebraic, generic) for $1\leq i\leq k$ then $s=k+1$ and the splitting points are $\rho_{1}=3$, $\rho_{i}=5+(i-1)$ for $i=2,\ldots,k+1$.
+Thus we are in the Case I.
+    Pay attention that, if $\alpha_{0}\in\mathbb{K}^{\prime}$ is not generic, the pictire can be somewaht different.
+For example, let $y=\sqrt{2}\tau^{6}+\tau^{7}$.
+Then the branch $C$ has the equation $f(x,y)=(y^{2}-2x^{3})^{2}-4\sqrt{2}yx^{5}-x^{7}=0$ and the conjugate of $C$ with respect to $G$, gCgC ($g\in G$ is given by $g(\sqrt{2})=-\sqrt{2}$), has the parametrization $(\tau^{4},-\sqrt{2}\tau^{6}+\tau^{7})$ and the equation $f^{\prime}(x,y)=gf(x,y)=(y^{2}-2x^{3})^{2}+4\sqrt{2}yx^{5}-x^{7}=0$.
+In this case the only splitting point is $\rho=5$ (also $\delta_{C}=5$).
+Notice that the intersection multiplicity of $C$ and $C^{\prime}=gC$ is $26$ and that the coordinate of the intersection point of the strict transform of $C$ (resp.
+$C^{\prime}$) is a certain $a\in\mathbb{Q}(\sqrt{2})\setminus\mathbb{Q}$ (resp.
+$g(a)$).
+  2. 2.
+    $\mathbb{K}^{\prime}$ is an infinite algebraic extension of $\mathbb{Q}=\mathbb{K}$.
+Then $s=\infty$ and we are in Case II.
+As in the previous case, in general, the splitting points are $3,6,7,\ldots$.
+  3. 3.
+    $\mathbb{K}^{\prime}$ is a trascendental extension of $\mathbb{Q}$.
+In this case we are in Case III.
+For simplicity, let us assume that $\alpha_{0}$ is trascendental over $\mathbb{Q}$.
+Then the process is interrumpted in the divisor $\delta_{\nu}=3$.
+In order to get a function $g\in{\mathcal{O}}_{\mathbb{C}^{2},0}$ such that $\nu(g)=2r+1\geq 7$ one needs that
+    $g=(y^{2}-\alpha_{0}^{2}x^{3})+\text{ higher order terms }\;.$
+    But in this case the equation can not be rational.
+So, in this case, the semigroup $S^{\mathbb{Q}}_{\nu}$ is $\langle 4,6\rangle\subset 2\mathbb{Z}$ and the valuation $\nu$ is $2$ times the divisorial valuation defined by the component $E_{3}$.
 Remark 4.
-Let $\mathbb{K}\subset\mathbb{C}$ (e.
-g., $\mathbb{K}=\mathbb{Q}$) and let $F\in\mathbb{K}[[x,y]]$ be irreducible.
+Let $F\in\mathbb{K}[[x,y]]$ be irreducible.
 The local ring of the irreducible curve (over $\mathbb{K}$) defined by $F$ is $R=\mathbb{K}[[x,y]]/(F)$.
-The integral closure $\overline{R}$ in its quotient field is the valuation ring $R_{\omega}$ of a discrete valuation $\omega$, its residual field $\mathbb{K}^{\prime}$ is a finite algebraic extension of $\mathbb{K}$ and could be realized and therefore identified with a subfield of $\mathbb{C}$.
-The equation $F$ defines a reduced algebroid curve over the complex field $\mathbb{C}$, each branch corresponds to an irreducible factor of $F$ in $\mathbb{C}[[x,y]]$.
+The integral closure $\overline{R}$ in its quotient field is the valuation ring $R_{\nu}$ of a discrete valuation $\nu$, its residual field $\mathbb{K}^{\prime}$ is a finite algebraic extension of $\mathbb{K}$ and could be realized and therefore identified with a subfield of $\mathbb{C}$.
+The equation $F=0$ defines a reduced algebroid curve over the complex field $\mathbb{C}$, each branch corresponds to an irreducible factor of $F$ in $\mathbb{C}[[x,y]]$.
 Let $C$ be one of these branches.
-The valuation $\nu=\nu_{C}$ defined by $C$ on $\mathbb{K}[[x,y]]$ coincides with the valuation $\omega$ (seen as a valuation on $\mathbb{K}[[x,y]]$ forgeting the quotient by $F$) and the Poincaré series (over $\mathbb{K}$) defined by $\nu$ is the Poincaré series defined by $\omega$.
-Note that all the branches are equisingular, the minimal resolution of $\{F=0\}$ (in the complex setting) coincides with the $\mathbb{K}$-resolution of $\nu_{C}$ and all the valuations $\nu_{C}$, for complex branches $C$ of $\{F=0\}$, restrict to the same valuation $\omega$ in $\mathbb{K}[[x,y]]$.
+The valuation $\nu_{C}$ defined by $C$ on $\mathbb{K}[[x,y]]$ coincides with the valuation $\nu$ (seen as a valuation on $\mathbb{K}[[x,y]]$ forgetting the quotient by $F$) and the Poincaré series (over $\mathbb{K}$) defined by $\nu_{C}$ is the Poincaré series defined by $\nu$.
+Note that all the branches are equisingular, the minimal resolution of $\{F=0\}$ (in the complex setting) coincides with the $G$-resolution of $\nu_{C}$ and all the valuations $\nu_{C}$, for complex branches $C$ of $\{F=0\}$, restrict to the same valuation $\omega$ in $\mathbb{K}[[x,y]]$.
 Moreover, if $G^{\prime}$ is the Galois group of the extension $\mathbb{K}\subset{\mathbb{L}}$, $\mathbb{L}$} the normal closure of $\mathbb{K}^{\prime}$ in $\mathbb{C}$, $G^{\prime}$ is the quotient of $G$ by the isotropy group of the field $\mathbb{L}$.
 This situation just corresponds to the case I described above.
+Let $C$ be a curve parametrized as in (2).
+The map $f\mapsto f(x(\tau),y(\tau))$ defines an homomorphism $\Phi:\mathbb{K}[[x,y]]\to\mathbb{C}[[\tau]]$.
+If $\ker(\Phi)\neq 0$, then one has $\ker(\Phi)=(F)$ for some irreducible $F\in\mathbb{K}[[x,y]]$ and we are in the above described situation, i.
+e., also in Case I.
+Otherwise, $\ker(\Phi)=(0)$ and we are in Case II or III.
+In this case the valuation $\nu_{C}$ is centered in the two-dimensional ring $\mathbb{K}[[x,y]]$.
+Case II corresponds to the case in which the residual field $\mathbb{K}_{\nu_{C}}$ of the valuation ring is an algebraic extension of $\mathbb{K}$ but not finite dimensional.
+Case III corresponds to the case in which $\mathbb{K}_{\nu_{C}}$ is a trascendental extension of $\mathbb{K}$.
 4 Subgroups of the Galois group, subfields, divisors and curvettes
 Further on we shall discuss only Cases I and II.
+We do not discuss Case III since in this case the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ is a multiple of a divisorial valuation whose Poincaré series is described in Section 7.
 Let $(E_{\sigma}\circ E_{\delta})$, $\sigma,\delta\in\Gamma$, be the intersection matrix of the components $E_{\sigma}$ (infinite one in Case II).
-The diagonal entry $E_{\sigma}\circ E_{\sigma}$ is the self-intersection number of the component $E_{\sigma}$ on the final modification in Case I and on a modification at a step, after which the points of $E_{\sigma}$ are not blown-up, in Case II.
+The diagonal entry $E_{\sigma}\circ E_{\sigma}$ is the self-intersection number of the component $E_{\sigma}$ on the final modification in Case I and, in Case II, a modification at a step, after which the points of $E_{\sigma}$ are not blown-up.
 One has $E_{\sigma}\circ E_{\sigma}<0$.
-The non-diagonal entry $E_{\sigma}\circ E_{\delta}$ is equal to 1 (respectively to 00) if these components intersect (do not intersect respectively).
+The non-diagonal entry $E_{\sigma}\circ E_{\delta}$ is equal to 1 (respectively to 0) if these components intersect (do not intersect respectively).
 Let $(m_{\sigma,\delta}):=-(E_{\sigma}\circ E_{\delta})^{-1}$.
-In Case II one can compute the entry $m_{\sigma,\delta}$ at a sufficiently late step (after which it stabilizes).
-In other words one can compute the intersection matrix and the inverse of it at each step and use their stabilization.
+In order not to treat infinite matrices, in Case II the entry $m_{\sigma,\delta}$ can be computed at a step of the resolution process after which the components $E_{\sigma}$ and $E_{\delta}$ are not modified.
 The number $m_{\sigma,\delta}$ can be interpreted as the intersection number of some germs of curves on $(\mathbb{C}^{2},0)$, so-called curvettes.
-A curvette correponding to the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a complex-analytic smooth curve $\gamma_{\sigma}$ (in Case II on a modification where $E_{\sigma}$ stabilizes) transversal to the component $E_{\sigma}$ at a smooth point of the exceptional divisor, that is not an intersection point with other components.
+A curvette corresponding to the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a complex-analytic smooth curve $\gamma_{\sigma}$ (in Case II on a modification after which the component $E_{\sigma}$ is not modified) transversal to the component $E_{\sigma}$ at a smooth point of the exceptional divisor, that is not an intersection point with other components.
 One can show that $m_{\sigma,\delta}$ is equal to the intersection number $C_{\sigma}\circ C_{\delta}$.
-For $\sigma\in\Gamma$, let $m_{\sigma}$ be the value of $\nu_{C}$ on the left-hand side of an equation $f_{\sigma}=0$ defining the curvette $C_{\sigma}$, i.
-e., the intersection number $C\circ C_{\sigma}$.
-One can see that, in Case I, $m_{\sigma}$ is equal to $m_{\sigma,\delta_{C}}$ for $\delta_{C}$ such that, on the surface of the resolution, the strict transform of the branch $C$ intersects the component $E_{\delta_{C}}$ and, in Case II, $m_{\sigma}$ is equal to $m_{\sigma,\delta}$ for $\delta$ on the subgraph of the usual resolution of the curve $C$ far enough from the initial component of the modification.
+For $\sigma\in\Gamma$, let $m_{\sigma}=\nu_{C}(f_{\sigma})$, where $f_{\sigma}=0$ is an equation defining a curvette $C_{\sigma}$.
+It is equal to the intersection number $C\circ C_{\sigma}$.
+One can see that, in Case I, $m_{\sigma}$ is equal to $m_{\sigma,\delta_{C}}$ for $\delta_{C}$ such that, on the surface of the resolution, the strict transform of the branch $C$ intersects the component $E_{\delta_{C}}$.
+In Case II, the vertex $\delta_{C}$ makes sense only at any step of the resolution process of the curve $C$.
+One can see that $m_{\sigma}=m_{\sigma,\delta}$ at a step of the resolution process when the branch $C$ itself is resolved (not in the $G$-invariant sense) and after which the component $E_{\sigma}$ is not modified.
 Isotropy subgroups and subfields.
-Let $G_{\sigma}$ be the isotropy subgroup of the component $E_{\sigma}$, i.
+Let, as above, $G_{\sigma}$ be the isotropy subgroup of the component $E_{\sigma}$, i.
 e., the isotropy subgroup of the vertex $\sigma$ in $\Gamma$.
@@
 For all vertices $\sigma\in\Gamma$ in between the vertex $\rho_{j}$ and $\rho_{j+1}$ ($\rho_{j}$ excluded, $\rho_{j+1}$ included) the isotropy subgroups $G_{\sigma}$ and the corresponding subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G_{j}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{j}$.
-In Case I, for all the vertices $\sigma\in\Gamma$ betwin $\rho_{s}$ and $\delta_{C}$ one has $G_{\sigma}=G_{s}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{s}$.
+In Case I, for all the vertices $\sigma\in\Gamma$ between $\rho_{s}$ and $\delta_{C}$ one has $G_{\sigma}=G_{s}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{s}$.
 For vertices from $G$-orbits of the discussed ones, the corresponding subgroups are conjugate to $G_{\sigma}$ and the subfields are obtained from $\mathbb{K}_{\sigma}$ by the shift by an element of the Galois group $G$.
 One has $G=G_{0}\supset G_{1}\supset\ldots\supset G_{s}$, $\mathbb{K}=\mathbb{K}_{0}\subset\mathbb{K}_{1}\subset\ldots\subset\mathbb{K}_{s}$.
-One has $[K_{j}:K_{j-1}]=[G_{j-1}:G_{j}]$.
+One has $\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]$.
 Standard charts.
@@
 On all the components $E_{\sigma}$ of the exceptional divisor except $E_{\rho_{j}}$ and those from its $G$-orbit, on each step the blow-up is made at one point with the coordinates $(0,a)$ in a standard chart with $a\in\mathbb{K}_{\sigma}$.
-In this case it can be convenient to change the coordinate $w$ by $\widetilde{w}=w/(w-a)$ or $\widetilde{w}=(w-a)/w$ so that the coordinates of the intersection points with the components on the geodesic between $\sigma_{0}$ and $\delta_{C}$ become equal to 00 and $\infty$.
+In this case it can be convenient to change the coordinate $w$ by $\widetilde{w}=w/(w-a)$ or $\widetilde{w}=(w-a)/w$ so that the coordinates of the intersection points with the components on the geodesic between $\sigma_{0}$ and $\delta_{C}$ become equal to 0 and $\infty$.
 On the component $E_{\rho_{j}}$, the blow-ups are made at a certain point $(0,a)$ with $a\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ (that is not from $\mathbb{K}_{\rho_{j}}=\mathbb{K}_{j-1}$) and at the points from its $G_{j-1}$-orbit.
-In this case a change as above is inappropriate.
+In this case a coordinate change as above cannot be used since it is not defined over $\mathbb{K}_{j-1}$.
 Special curvettes.
@@
 In other words, the curve $\gamma_{\sigma}$ is invariant under the action of $G_{\sigma}$.
-In particular one can take the curve $\{u=\tau$, $w=c=\text{const}\}$ in a standart chart with $c\in\mathbb{K}_{\sigma}$ ($u=0$ is the equation of $E_{\sigma}$ in the chart).
+In particular one can take the curve $\{u=\tau$, $w=c=\text{const}\}$ in a standard chart with $c\in\mathbb{K}_{\sigma}$ ($u=0$ is the equation of $E_{\sigma}$ in the chart).
 Definition 3.
@@
 5 The semigroup Poincaré series of a curve valuation
-Let $\nu(=\nu_{C})$ be the valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ defined by a curve $C\subset(\mathbb{C}^{2},0)$ and assume that $\nu_{C}$ is a valuation of the type I or II described in Section 3.
+Let $\nu(=\nu_{C})$ be the valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ defined by an irreducible curve $C\subset(\mathbb{C}^{2},0)$ and assume that $\nu_{C}$ is a valuation of the type I or II described in Section 3.
 Let $P^{S}_{C}(t)=\sum_{m\in S^{\mathbb{K}}_{\nu}}t^{m}$ be the semigroup Poincaré series of $\nu_{C}$ (see Definition 2).
-For $\sigma\in\check{\Gamma}$, let $M_{\sigma}$ be the value of $\nu$ on the left-hand side of the equation $F_{\sigma}=0$ definig a $G_{\sigma}$-curvette at the component $E_{\sigma}$.
+For $\sigma\in\check{\Gamma}$, let $M_{\sigma}=\nu(F_{\sigma})$ where $F_{\sigma}=0$ is the equation defining a $G_{\sigma}$-curvette at the component $E_{\sigma}$.
 One can see that $M_{\sigma}=\sum_{[g]\in G/G_{\sigma}}m_{g\sigma}$.
@@
 $P^{S}_{C}(t)=\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\;.$
-Let $S_{C}$ be the usual semigroup of the complex curve $C$, i.
-e., $S_{C}=\{\nu(f):f\in{\mathcal{O}}_{\mathbb{C}^{2},0}\}$.
-The set of multiplicities $\{m_{\sigma_{0}},\ldots,m_{\sigma_{g}}\}$ is the minimal set of generators of $S_{C}$.
+Let $S^{\mathbb{C}}_{C}$ be the usual semigroup of the complex branch $C$, i.
+e., $S^{\mathbb{C}}_{C}=\{\nu(f):f\in{\mathcal{O}}_{\mathbb{C}^{2},0}\}$.
+The set of multiplicities $\{m_{\sigma_{0}},\ldots,m_{\sigma_{g}}\}$ is the minimal set of generators of $S^{\mathbb{C}}_{C}$.
 Let $e_{i}:=\gcd\{m_{\sigma_{0}},\,\ldots,m_{\sigma_{i}}\}$ for $i=0,1,\ldots,g$ and let $N_{i}=e_{i-1}/e_{i}$ for $i=1,2,\ldots,g$.
@@
 Notice that $e_{\sigma}(f)$ coincides with the intersection multiplicity of the strict transform of $\{f=0\}$ and $E_{\sigma}$ in the surface $X_{\sigma}$.
-For $j=1,\ldots,s$, let $\ell_{j}:=[\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]:=|\frac{G_{j-1}}{G_{j}}|$.
+For $j=1,\ldots,s$, let $\ell_{j}:=[\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]=\#(G_{j-1}/{G_{j}})$.
 Lemma 1.
-Let $\omega\in\check{\Gamma}$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$ and $\tau_{p}\leq\omega<\tau_{{p}+1}$. One has
+Let $\omega\in\check{\Gamma}$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$ and $\tau_{p}<\omega\leq\tau_{{p}+1}$. One has
 $M_{\omega}=\sum_{i=0}^{{p}}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
-for some non negative integers $k_{i}$ $(0\leq i\leq{p})$ and positive integers $r_{j}$ $(1\leq j\leq d)$.
+for some non negative integers $k_{i}$ $(0\leq i\leq{p})$, $k_{i}<N_{i}$ for $i\geq 1$, and positive integers $r_{j}$ $(1\leq j\leq d)$.
 Proof..
@@
 Figure 6: The position of the vertices.
-Moreover, it is known that $m_{\omega}\in\langle m_{0},\,\ldots,m_{p}\rangle$ and so $m_{\omega}=k_{0}m_{0}+\sum_{i=1}^{p}k_{i}m_{i}$, for some $k_{0}\geq 0$ and $0\leq k_{i}<N_{i}$ for $i\geq 1$ (which are unique with this restrictions).
-Then one can write
+Moreover, it is known that $m_{\omega}\in\langle m_{0},\,\ldots,m_{p}\rangle$, where $\langle m_{0},\,\ldots,m_{p}\rangle$ stands for the subsemigroup of $\mathbb{Z}_{\geq 0}$ generated by $m_{0},\,\ldots,m_{p}$, i.e.
+$\langle m_{0},\,\ldots,m_{p}\rangle:=\{n\in\mathbb{N}:n=a_{0}m_{0}+\cdots+a_{p}m_{p}\text{ with }a_{i}\in\mathbb{Z}_{\geq 0}\}\;,$
+and so $m_{\omega}=k_{0}m_{0}+\sum_{i=1}^{p}k_{i}m_{i}$, for some $k_{0}\geq 0$ and $0\leq k_{i}<N_{i}$ for $i\geq 1$ (which are unique with these restrictions). Then one can write
 $M_{\omega}=\sum_{0}^{p}k_{i}m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$ (6)
-Let us assume that $q<{p}$.
+Let us assume that q<pq<{p}, note that in this case we have $\tau_{q-1}<\rho_{d}\leq\tau_{q}<\tau_{p}\leq\omega$ (see Figure 6).
 For $j\leq d$ one has $e_{\rho_{j}}(\varphi_{\omega})/e_{\tau_{{p}-1}}(\varphi_{\omega})=e_{\rho_{j}}(\varphi_{{p}})/e_{\tau_{{p}-1}}(\varphi_{p})$, moreover $e_{\tau_{{p}-1}}(\varphi_{p})=1$, so $e_{\rho_{j}}(\varphi_{\omega})=e_{\tau_{{p}-1}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})$.
-It is known that $e_{i}=e_{\tau_{i}}(\varphi_{C})$, $i\leq g$ and $\varphi_{C}$ is an equation of the complex curve $C$.
+It is known that $e_{i}=e_{\tau_{i}}(\varphi_{C})$, $i\leq g$, where $\varphi_{C}=0$ is an equation of the complex curve $C$.
 As a consequence, for $i\leq p$, one has $N_{i}=e_{i-1}/e_{i}=e_{\tau_{i-1}}(\varphi_{\omega})/e_{\tau_{i}}(\varphi_{\omega})$ and so
-$e_{\rho_{j}}(\varphi_{\omega})=N_{p}e_{\tau_{p}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})\;.$
-The same computations applied to $\varphi_{i}$ (instead of $\varphi_{\omega}$) and $\varphi_{i-1}$ (for $\tau_{i}>\tau_{i-1}>\rho_{j}$) show that $e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\tau_{i-1}}(\varphi_{i})e_{\rho_{j}}(\varphi_{i-1})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$.
-Using this equality in the rigth part of the equation (6) and also the expression of $M_{i}$ above, one gets:
+$e_{\rho_{j}}(\varphi_{\omega})=N_{p}e_{\tau_{p}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})\;.$ (7)
+For $i,j$ such that $\tau_{i}>\tau_{i-1}>\rho_{j}$, the same computations applied to $\varphi_{i}$ (instead of $\varphi_{\omega}$) and $\varphi_{i-1}$ (instead of $\varphi_{p}$) show that
+$e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\tau_{i-1}}(\varphi_{i})e_{\rho_{j}}(\varphi_{i-1})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})\;.$ (8)
+Using equality (7) in the right part of equation (6), one gets:
 $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{\omega})m_{\rho_{j}}$ $=N_{p}e_{\tau_{p}}(\varphi_{\omega})\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}=$
 $=k_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}+k^{\prime}_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}$
-with $k^{\prime}_{p}=N_{p}e_{\tau_{p}}(\varphi_{\omega})-k_{p}\geq 1$, and so
+with $k^{\prime}_{p}=N_{p}e_{\tau_{p}}(\varphi_{\omega})-k_{p}\geq N_{p}-k_{p}\geq 1$, with $k_{p}$ being the integer appearing in equation (6).
+Now, using the expression of $M_{i}$ in (5) for $i=p$, one has
 $M_{\omega}=\sum_{0}^{{p}-1}k_{i}m_{i}+k_{p}M_{p}+k^{\prime}_{p}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{p})m_{\rho_{j}}\;.$
-In the same way, using repeatedly that $e_{\rho_{i}}(\varphi_{j})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$ for $i\geq q$ one gets:
-$M_{\omega}=(\sum_{0}^{q-1}k_{i}m_{i})+\sum_{i=q}^{p}k_{i}M_{i}+r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}$ (7)
-for some integer $r_{d}\geq 1$. Now, using that $L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}$ for $j\leq d-1$ in the last summand of the above equality (7), one has
-$r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=r_{d}[G_{d-1}:G_{d}]\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}+r_{d}L^{d}_{d}e_{\rho_{d}}(\varphi_{q})m_{\rho_{d}}\;.$
-Using that $L^{d}_{d}=[G_{d-1}:G_{d}]-1$, $e_{\rho_{j}}(\varphi_{q})=e_{\rho_{j}}(\varphi_{\rho_{d}})$, $e_{\rho_{d}}(\varphi_{\rho_{d}})=1$ and the corresponding expression for $M_{\rho_{d}}$ in the same terms as at the beginning of the proof for $M_{\omega}$, the last equality can be written as:
+In the same way, using repeatedly that $e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$ and the equation (5) for $M_{i}$, $i\geq q$, one gets:
+$M_{\omega}=\sum_{i=0}^{q-1}k_{i}m_{i}+\sum_{i=q}^{p}k_{i}M_{i}+r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}$ (9)
+for some integer $r_{d}\geq 1$. Now, one can use the equality $L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}$ for $j\leq d-1$ and (8) in the last summand of (9) to get:
+$r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=r_{d}[G_{d-1}:G_{d}]\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}+r_{d}L^{d}_{d}e_{\rho_{d}}(\varphi_{q})m_{\rho_{d}}\;.$ (10)
+Using that $L^{d}_{d}=[G_{d-1}:G_{d}]-1$, $e_{\rho_{j}}(\varphi_{q})=e_{\rho_{j}}(\varphi_{\rho_{d}})$, $e_{\rho_{d}}(\varphi_{\rho_{d}})=1$ and the corresponding expression for $M_{\rho_{d}}$ in the same terms as we use at the beginning of the proof for $M_{\omega}$, the equality (10) can be written as:
 $r_{d}([G_{d-1}:G_{d}]-1)\left(\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{\rho_{d}})m_{\rho_{j}}+m_{\rho_{d}}\right)+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=$
 $=r_{d}([G_{d-1}:G_{d}]-1)M_{\rho_{d}}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\;.$
-Thus, taking into account that $\ell_{j}=[G_{j-1}:G_{j}]$ and the last equality, the equation (7) for $M_{\omega}$ becomes:
+Thus, using this equality and taking into account that $\ell_{j}=[G_{j-1}:G_{j}]$, the equation (9) for $M_{\omega}$ becomes:
 $M_{\omega}=\sum_{i=q}^{p}k_{i}M_{i}+r_{d}(\ell_{d}-1)M_{\rho_{d}}+\left(\sum_{0}^{q-1}k_{i}m_{i}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\right)\;.$ (11)
-If $d=1$, one has $m_{i}=M_{i}$ for $i=0,\ldots,q-1$ and the proof is finished.
-Otherwise, we apply the induction hypothesis to the last two summands in (8) (in brackets).
-Notice that the factor $r_{d}\geq 1$ does not affects the proof above and the terms $e_{\rho_{j}}(\varphi_{q})$ (corresponding to a curvette at $E_{\sigma_{q}}$) could be treated (arithmetically) as we did with $e_{\rho_{j}}(\varphi_{\omega})$ for $j<d$.
-As a consequence, we have:
+If $d=1$, one has $m_{i}=M_{i}$ for $i=0,\ldots,q-1$ and the proof is finished. Otherwise, we repeat the above computations for the last two summands in (11) (in brackets):
+$\sum_{0}^{q-1}k_{i}m_{i}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\;.$ (12)
+To do that, note that the equation (5) remains valid for $M_{i}$, $i\leq q$, and that the equation (12) is essentially the same as the one in the right hand side of (6):
+  1. 1.
+    The factor $r_{d}\geq 1$ does not affect the computations in the proof.
+  2. 2.
+    Equation (8) could be applied to $\varphi_{p}$ (and to the succesive $\varphi_{i}$, $i<p$) instead of equation (7) for $\varphi_{\omega}$.
+  3. 3.
+    The succesive integers $k^{\prime}_{i}\geq N_{i}-k_{i}$ are always greater or equal to 1.
+As a consequence, we get:
 $M_{\omega}$ $=\sum_{i:\rho_{1}<\tau_{i}\leq\tau_{p}}k_{i}M_{i}+\sum_{i:\tau_{i}\leq\rho_{1}}k_{i}m_{i}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
@@
 for some integers $r_{j}\geq 1$ ($1\leq j\leq d$). ∎
-Remark 5.
-For $\sigma\in\check{\Gamma}$, $\tau_{k}\leq\sigma<\tau_{k+1}$ it is known that $m_{\sigma}\in\langle m_{\sigma_{0}},\ldots,m_{\sigma_{k}}\rangle$ or equivalently that $e_{k}$ divides $m_{\sigma}$ (see [12].
-Taken into account the expression for $M_{i}$ given in equation (5), this implies that $e_{i-1}$ divides $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{i}}$ and so $\gcd(M_{0},\,\ldots,M_{i})=\gcd(m_{0},\,\ldots,m_{i})=e_{i}$ and
-$N_{i}=e_{i-1}/e_{i}=\gcd(M_{0},\,\ldots,M_{i-1})/\gcd(M_{0},\,\ldots,M_{i})\,.$
-Lemma 2.
-For $i=1,\ldots,g$ one has $(N_{i}-1)M_{\sigma_{i}}\notin\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$, $N_{i}M_{\sigma_{i}}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$ and $N_{i-1}M_{\sigma_{i-1}}<M_{\sigma_{i}}$.
-In particular, $\{M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ and, moreover, $S$ is the semigroup of values of a complex analytic irreducible curve.
+Proposition 1.
+The set of integers $\{M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}\}$ is a set of generators of the semigroup of values $S^{\mathbb{K}}_{C}$ of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
 Proof..
-Note that, as a consequence of the Remark 5, $(N_{i}-1)M_{i}\notin\langle M_{0},\ldots,M_{i-1}\rangle$.
-Let us prove that for $i=1,\ldots,g$, one has that $N_{i-1}M_{i-1}<M_{i}$.
-If $\rho_{d}\leq\tau_{i-1}<\tau_{i}$, the equation (5) for $M_{i}$ together with the equality $e_{\rho_{i}}=N_{i-1}e_{\rho_{i}}(\varphi_{i-1})$ for $i\leq d$ implie that
-$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}=m_{i}+N_{i-1}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}>$
-$>N_{i-1}\left(m_{i-1}+\sum_{i=j}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
-If $\tau_{i-1}<\rho_{i}\leq\tau_{i}$ the equation (5) for $M_{i-1}$ is $M_{i-1}=m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}$ for some $d^{\prime}<d$.
-Thus, using the equality
-$L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}=\ell_{d-1}L^{d-1}_{j}=\ell_{d-1}\cdots\ell_{d^{\prime}}L^{d^{\prime}}_{j}$
-one gets
-$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)$
-$>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
-It is known that the proved properties implies that the semigroup $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ is the semigroup of a complex irreducible plane curve (see [1]).
-Proposition 1.
-The set of integers $\{M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of the semigroup of values $S^{\mathbb{K}}_{C}$ of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
-Proof..
-By Lemma 2 it is enough to prove that $S^{\mathbb{K}}_{\nu}$ is generated by $M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}$.
 It is clear that the semigroup $S^{\mathbb{K}}_{\nu}$ is generated by the integers $M_{\sigma}$, for all $\sigma\in\check{\Gamma}$.
 Thus one has to show that $M_{\sigma}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ for any $\sigma$.
-Let $[\alpha,\beta]$ denote the geodesic in $\check{\Gamma}$ between the vertices $\alpha$ and $\beta$. If $\omega\in[\sigma_{0},\delta_{C}]$, then Lemma 1 gives
-$M_{\omega}=\sum_{i=0}^{h}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}$
-where $\rho_{d}<\omega\leq\rho_{d+1}$, $\tau_{h}\leq\omega<\tau_{h+1}$ and $r_{j},k_{i}$ are non-negative integers.
+Let $\alpha,\beta]$ denote the geodesic in $\check{\Gamma}$ between the vertices $\alpha$ and $\beta$. If $\omega\in[\sigma_{0},\delta_{C}]$, Lemma 1 gives
+$M_{\omega}=\sum_{i=0}^{h}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}\;,$
+where $\rho_{d}<\omega\leq\rho_{d+1}$, $\tau_{h}<\omega\leq\tau_{h+1}$, $k_{i}$ are non-negative integers and $r_{j}$ are positive integers.
 The same equation applied to $\rho_{j}$ gives us that $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{i}\rangle$ if $\tau_{i}\leq\rho_{j}<\tau_{i+1}$.
@@
 Note that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$.
+Remark 7.
+For $\sigma\in\check{\Gamma}$, $\tau_{k}<\sigma\leq\tau_{k+1}$, it is known that $m_{\sigma}\in\langle m_{\sigma_{0}},\ldots,m_{\sigma_{k}}\rangle$ or equivalently that $e_{k}$ divides $m_{\sigma}$ (see [15]).
+Taken into account the expression for $M_{i}$ given in equation (5), this implies that $e_{i-1}$ divides $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{i}}$ and so $\gcd(M_{0},\,\ldots,M_{i})=\gcd(m_{0},\,\ldots,m_{i})=e_{i}$ and
+$N_{i}=e_{i-1}/e_{i}=\gcd(M_{0},\,\ldots,M_{i-1})/\gcd(M_{0},\,\ldots,M_{i})\,.$
+Proposition 2.
+For $i=1,\ldots,g$, one has $(N_{i}-1)M_{\sigma_{i}}\notin\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$, $N_{i}M_{\sigma_{i}}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$ and $N_{i-1}M_{\sigma_{i-1}}<M_{\sigma_{i}}$.
+In particular, $\{M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of the semigroup $S^{\mathbb{K}}_{C}:=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ and, moreover, $S^{\mathbb{K}}_{C}$ is the semigroup of values of a complex analytic irreducible curve.
 Proof..
-(Of Theorem 1) It follows directly from Proposition 1 and Lemma 2 taking into account that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$ for all $i\geq 1$. ∎
+If $(N_{i}-1)M_{i}\in\langle M_{0},\,\ldots,M_{i-1}\rangle$ then, as a consequence of Remark 7, $e_{i-1}$ divides $(N_{i}-1)M_{i}$ and so $e_{i-1}$ divides $(N_{i}-1)m_{i}$.
+But this contradicts the properties of the minimal set of generators of the semigroup $S^{\mathbb{C}}_{C}$.
+Note that $N_{i}M_{i}=M_{\tau_{i}}$, then $N_{i}M_{i}\in\langle M_{0},\,\ldots,M_{i-1}\rangle$ as a consequence of Proposition 1 (more precisely of the first part of its proof).
+Let us prove that, for $i=1,\ldots,g$, one has that $N_{i-1}M_{i-1}<M_{i}$.
+If $\rho_{d}\leq\tau_{i-1}<\tau_{i}$, the equation (5) for $M_{i}$ together with the equality $e_{\rho_{i}}=N_{i-1}e_{\rho_{i}}(\varphi_{i-1})$ for $i\leq d$ imply that
+$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}=m_{i}+N_{i-1}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}>$
+$>N_{i-1}\left(m_{i-1}+\sum_{i=j}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
+If $\tau_{i-1}<\rho_{i}\leq\tau_{i}$ the equation (5) for $M_{i-1}$ is $M_{i-1}=m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}$ for some $d^{\prime}<d$.
+Thus, using the equality
+$L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}=\ell_{d-1}L^{d-1}_{j}=\ell_{d-1}\cdots\ell_{d^{\prime}}L^{d^{\prime}}_{j}\;,$
+one gets
+$M_{i}$ $=m_{i}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{j}}>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)$
+$>N_{i-1}\left(m_{i-1}+\sum_{j=1}^{d^{\prime}}L^{d^{\prime}}_{j}e_{\rho_{j}}(\varphi_{i-1})m_{\rho_{j}}\right)=N_{i-1}M_{i-1}\;.$
+It is known that the proved properties implies that the semigroup $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ is the semigroup of a complex irreducible plane curve (see [1]).
+Proof of Theorem 1. The statement follows directly from Propositions 1 and 2 taking into account that $M_{\tau_{i}}=N_{i}M_{\sigma_{i}}$ for all $i\geq 1$. $\Box$
 Remark 8.
-As we have proved, $S^{\mathbb{K}}_{\nu}$ is the semigroup of a complex analytic branch $C^{\prime}$ (and so $P^{S}_{C}(t)$ is the usual Poincaré serie of it) different from $C$ if $m_{\rho_{1}}<m_{\tau_{g}}$.
+As we have proved, $S^{\mathbb{K}}_{\nu}$ is the semigroup of a complex analytic branch $C^{\prime}$ (and so $P^{S}_{C}(t)$ is the usual Poincaré series of it) different from $C$ if $m_{\rho_{1}}<m_{\tau_{g}}$.
 We do not know a geometric reason for the existence of the curve $C^{\prime}$.
-6 The classical Poincaré series
+6 The classical Poincaré series of a curve valuation
 Theorem 2.
 In Cases I and II, one has
-$P_{C}(t)$ $=$ $P_{C}^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
-$=$ $\frac{\prod_{i=0}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
+$P^{\mathbb{K}}_{C}(t)$ $=$ $P_{C}^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
+$=$ $\frac{\prod_{i=1}^{g}(1-t^{M_{\tau_{i}}})}{\prod_{i=0}^{g}(1-t^{M_{\sigma_{i}}})}\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}$
 ($s=\infty$ in Case II).
-Proof..
-Let $P_{C}(t)=\sum_{v=0}^{\infty}a_{v}t^{v}$.
+Before proving the Theorem, we elaborate some statements.
+Let $P^{\mathbb{K}}_{C}(t)=\sum_{v=0}^{\infty}a_{v}t^{v}$.
 For a collection $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$, $k_{\sigma}\in\mathbb{Z}_{\geq 0}$, let ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ be the set of germs $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that the intersection multiplicity of the strict transform of the zero-level set $\{f=0\}$ with the component $E_{\sigma}$ of the exceptional divisor $D$ is equal to $k_{\sigma}$ and, moreover, this strict transform intersects $D$ only at smooth points of the total transform $\pi^{-1}(C)$ of the curve $C$.
@@
 Remark 9.
-Without loss of generality, we shall use the assumption that, for a fixed $v=\nu(\{k_{\sigma}\})$, the latter condition holds automatically.
+Without loss of generality, we shall use the assumption that, for a fixed $v=\nu(\{k_{\sigma}\})$, the condition that the zero-level set $\{f=0\}$ with $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ intersects $D$ only at smooth points holds automatically.
 This can be achieved if one makes sufficiently many additional blow-ups at intersection points of the components of the total transform $\pi^{-1}(C)$.
-Let $F^{\{k_{\sigma}\}}$ be the image of ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ in the quotient $J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)\cong\mathbb{C}$ with $v=\nu(\{k_{\sigma}\})$.
-Insted of writing that $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ (and explaining which $\{k_{\sigma}\}$ is considered), we shall sometimes write that $f$ corresponds to the representation $v=\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$ or simply to the expression $\sum_{\sigma\in\check{\Gamma}}k_{\sigma}M_{\sigma}$.
+Let $F^{\{k_{\sigma}\}}$ be the image of ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ in the quotient $J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)\cong\mathbb{C}$ with $v=\nu(\{k_{\sigma}\})$.(Of course $F^{\{k_{\sigma}\}}\subset J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\subset J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)$.) Let $\overline{F}^{\{k_{\sigma}\}}$ be the linear span over $\mathbb{K}$ of $F^{\{k_{\sigma}\}}$.
 One has:
   1. 1)
-    $\bigcup_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}F^{\{k_{\sigma}\}}=J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1);$
+    $\bigcup_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}F^{\{k_{\sigma}\}}=J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\setminus\{0\};$
   2. 2)
-    for each collection $\{k_{\sigma}\}$, $F^{\{k_{\sigma}\}}$ is the complement to a (finite) arrangement of (proper) vector $\mathbb{K}$-subspaces in the vector $\mathbb{K}$-space $\overline{F}^{\{k_{\sigma}\}}$ (the linear span of $F^{\{k_{\sigma}\}}$).
-Let $d^{\{k_{\sigma}\}}$ be the dimension (over $\mathbb{K}$) of the vector space $\overline{F}^{\{k_{\sigma}\}}$. One obviously has
-$a_{v}=\max_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
-For $j=1,2,\ldots,s+1$ ($s$ may be equal to +$\infty$+ ∞), let $A_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic $[\rho_{j},\delta_{C}]$ from $\rho_{j}$ to $\delta_{C}$ (including the ends).
+    for each collection $\{k_{\sigma}\}$, $F^{\{k_{\sigma}\}}$ is the complement to a finite hyperplane arrangement in the vector $\mathbb{K}$-space $\overline{F}^{\{k_{\sigma}\}}$.
+Let $d^{\{k_{\sigma}\}}$ be the dimension (over $\mathbb{K}$) of the vector space $\overline{F}^{\{k_{\sigma}\}}$.
+One obviously has $a_{v}=0$ if there are no $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$,
+$a_{v}=\max_{\{k_{\sigma}\}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}{\text{\ \ \ otherwise}}\,.$
+For $j=1,2,\ldots,s+1$ ($s$ may be equal to $\infty$), let $A_{j}$ be the set of collections $\{k_{\sigma}\}$ such that $k_{\sigma}=0$ for all $\sigma$ on the geodesic $\rho_{j},\delta_{C}]$ from $\rho_{j}$ to $\delta_{C}$ (including the ends).
 For $j=s+1$ (if $s<+\infty$), we assume this geodesic to be empty.
@@
 $a^{(j)}_{v}=\max_{\{k_{\delta}\}\in A_{j}:\nu(\{k_{\sigma}\})=v}d^{\{k_{\sigma}\}}.$
-One obviously has $P_{C}(t)=P^{(s+1)}_{C}(t)$ if $s<+\infty$ and, if $s=+\infty$ (i.
-e., if the process of the $\mathbb{K}$-resolution of the curve $C$ is infinite), $P_{C}(t)$ is, in a natural sense, the limit of $P^{(j)}_{C}(t)$ for $j\to+\infty$.
+One obviously has $P^{\mathbb{K}}_{C}(t)=P^{(s+1)}_{C}(t)$ if $s<+\infty$ and, if $s=+\infty$ (i.
+e., in Case II : the process of the $G$-resolution of the curve $C$ is infinite), $P^{\mathbb{K}}_{C}(t)$ is the limit of $P^{(j)}_{C}(t)$ for $j\to+\infty$ in “the $\mathfrak{m}$-adic topology” ($\mathfrak{m}$ is the maximal ideal in $\mathbb{Z}[t]$).
+The latter means that, for any $v$, $a_{v}=a_{v}^{(j)}$ for $j$ large enough.
 For a function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}$, let $\widetilde{f}=f\circ\pi$ be the lifting of $f$ to the surface of the resolution.
-In order to compute (or evaluate) the dimension $d^{\{k_{\sigma}\}}$, we shall regularly use the following construction.
+In order to compute (or evaluate) the dimension $d^{\{k_{\sigma}\}}$, we shall regularly use the following construction exploited first in [5].
 Let $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ be a fixed function with $\nu(h)=\nu(\{k_{\sigma}\})$.
 (In fact we shall take a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.) For each function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ we shall consider the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$.
-The function $\Psi$ is regular and different from zero at the intersection point of the exceptional divisor with the strict transform of the curve $C$.
-The set of values of these ratios for all $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ is isomorphic to the set $F^{\{k_{\sigma}\}}$.
+Let $q$ of the exceptional divisor with the strict transform of the curve $C$.
+The function $\Psi$ is regular in a neighbourhood of $q$ and its value $\Psi(q)$ is different from zero.
+The set $\left\{({\widetilde{f}}/{\widetilde{h}})(q):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}\right\}$ of values of these ratios at the point $q$ as a subset of $\mathbb{C}$ is obtained from $F^{\{k_{\sigma}\}}$, also considered as a subset of $\mathbb{C}\cong J^{\mathbb{C}}(v)/J^{\mathbb{C}}(v+1)$ with $v=\nu(\{k_{\sigma}\})$, by multiplication by a non-zero factor.
 This follows from the following arguments.
 The set $F^{\{k_{\sigma}\}}$ is just the set of the coefficients $a(f)$ in Equation (3) for $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.
-The coefficient $a(f)$ is equal to the coefficient $a(h)$ (a fixed number) mutiplied by the value of the ratio $\Psi$.
-If $\rho_{1}\neq\sigma_{0}$, the statement follows from the following six lemmas (and two Propositions). The necessary changes in the case $\rho_{1}=\sigma_{0}$ will be discussed later.
+The coefficient $a(f)$ is equal to the coefficient $a(h)$ (a fixed number) multiplied by the value of the ratio $\Psi$.
+Assume first that $\rho_{1}\neq\sigma_{0}$. The necessary changes in the case $\rho_{1}=\sigma_{0}$ will be discussed later.
 Lemma 2.
@@
 Any $v\in S^{\mathbb{K}}_{C}$ can be represented in the form $v=\sum_{i=0}^{g}k_{i}M_{\sigma_{i}}$ (Proposition 1).
-Any function $f$ corresponding to this representation gives a non-zero element in the corresponding $F^{\{k_{\sigma}\}}$.
+Any function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ with $k_{\sigma_{i}}=k_{i}$, $k_{\sigma}=0$ otherwise gives a non-zero element in the corresponding $F^{\{k_{\sigma}\}}$.
 Therefore $d^{\{k_{\sigma}\}}\geq 1$ and $a^{(1)}_{v}\geq 1$.
@@
 Let $\{k_{\sigma}\}\in A_{1}$ and let $h$ be a fixed function from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$.
-For $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{1}}$ and $E_{g\delta_{C}}$ for all $g\in G=G_{0}$.
+For $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ has no zeroes or poles on all the components from the union of the geodesics $\rho_{1},g\delta_{C}]$ for all $g\in G=G_{0}$.
 Therefore this is a (non-zero) constant on them.
@@
 ∎
-Lemmas 3 and 4 obviously imply the following statement.
+Lemmas 2 and 3 obviously imply the following statement.
 Proposition 3.
@@
 The proof is essentially the same as the one of Lemma 3.
-For $\{k_{\sigma}\}\in A_{j}$ and a fixed function $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ ($f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$) has no zeroes or poles on all the components from the union of the geodesics between $E_{\rho_{j}}$ and $E_{g\delta_{C}}$ for all $g\in G_{j-1}$.
+For $\{k_{\sigma}\}\in A_{j}$ and a fixed function $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$, the ratio $\Psi={\widetilde{f}}/{\widetilde{h}}$ ($f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$) has no zeroes or poles on all the components from the union of the geodesics $\rho_{j},g\delta_{C}]$ for all $g\in G_{j-1}$.
 Therefore this is a (non-zero) constant on them.
@@
 Lemma 5.
-Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation, assume that $\ell\leq\ell_{i}-1$.
+Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$ and such that $b-M_{\rho_{j}}\notin S_{C}^{\mathbb{K}}$.
+Assume that $\ell\leq\ell_{j}-1$ $(\ell_{j}=[\mathbb{K}_{j}:\mathbb{K}_{j-1}])$.
 We also suppose that $v$ is not of the form $M_{\varkappa}+b^{\prime}$ with $\varkappa>\rho_{j}$ and $b^{\prime}\in S^{\mathbb{K}}_{C}$.
-(The corresponding case will be anayzed in Lemma 8.) Then $a_{v}^{(j+1)}=\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b}^{(j)}$.
+(The corresponding case will be analyzed in Lemma 7.) Then $a_{v}^{(j+1)}=\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b}^{(j)}$.
 Proof..
@@
 One has $\mathbb{K}_{j}=\mathbb{K}_{j-1}[z_{0}]$ and $1$, $z_{0}$, $z_{0}^{2}$, …, $z_{0}^{\ell_{j}-1}$ is a basis of the vector space $\mathbb{K}_{j}$ over $\mathbb{K}_{j-1}$.
-Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\rho_{j}}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\rho_{j}$, $k^{2}_{\rho_{j}}=0$.
+Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $\{k_{\sigma}\}=\{k^{1}_{\sigma}\}+\{k^{2}_{\sigma}\}$, where $k^{1}_{\rho_{j}}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\rho_{j}$, $k^{2}_{\rho_{j}}=0$.
 A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$.
 Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$.
-For certainty, as $h_{1}$ we shall take the $\ell^{\prime}$-th power of the left hand side of the equation of the $G_{j-1}$-curvette at the point $w=1$ of the component $E_{\rho_{j}}$.
+For certainty, as $h_{1}$ we shall take the $\ell^{\prime}$-th power of the left hand side of the equation of a $G_{j-1}$-curvette at the point $w=1$ of the component $E_{\rho_{j}}$.
 Let $h=h_{1}h_{2}$.
 One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$.
-The function $\Psi_{2}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesics between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j-1}$ (including both end points of the geodesic).
+The function $\Psi_{2}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesics between $\rho_{j}$ and $g\delta_{C}$ with $g\in G_{j-1}$ (including the both end points of the geodesic).
 Its value on them belongs to $\mathbb{K}_{j-1}$ and the maximal possible dimension (over $\mathbb{K}$) of the set of its values is $a_{b}^{(j)}$.
@@
 Moreover, its value at $w=\infty$, that is $c$, belongs to $\mathbb{K}$.
-(This follows from the fact that this ratio is constant on the union of the components from the geodesic $[\sigma_{0},\rho_{j})$.) Thus up to the costant $(z_{0}-1)^{\ell^{\prime}}$ (which does not influence the dimensions) the value of $\Psi$ on the union of the components from the geodesic $(\rho_{j},\delta_{C}]$ is constant and is equal to $c\prod_{p=1}^{\ell^{\prime}}(z_{0}-w_{p})=c(z_{0}^{\ell^{\prime}}-(w_{1}+\ldots+w_{\ell^{\prime}})z_{0}^{\ell^{\prime}-1}+\ldots\pm(w_{1}\cdot\ldots\cdot w_{\ell^{\prime}}))$.
+(This follows from the fact that this ratio is constant on the union of the components from the geodesic $\sigma_{0},\rho_{j})$.) Thus up to the constant $(z_{0}-1)^{\ell^{\prime}}$ (which does not influence the dimensions) the value of $\Psi_{1}$ on the union of the components from the geodesic $(\rho_{j},\delta_{C}]$ is constant and is equal to $c\prod_{p=1}^{\ell^{\prime}}(z_{0}-w_{p})=c(z_{0}^{\ell^{\prime}}-(w_{1}+\ldots+w_{\ell^{\prime}})z_{0}^{\ell^{\prime}-1}+\ldots\pm(w_{1}\cdot\ldots\cdot w_{\ell^{\prime}}))$.
 Therefore the value of the ratio $\Psi$ on this union is equal to $c(\alpha_{\ell^{\prime}}z_{0}^{\ell^{\prime}}+\alpha_{\ell^{\prime}-1}z_{0}^{\ell^{\prime}-1}+\alpha_{1}z_{0}+\alpha_{0})$, where $\alpha_{0}$, …, $\alpha_{\ell^{\prime}-1}$ are generic points of $\mathbb{K}_{j-1}$ (because we can choose arbitrarily $w_{1},\,\ldots,w_{\ell^{\prime}}$ in $\mathbb{K}_{j-1}$) and $\alpha_{\ell^{\prime}}$ (also a point of $\mathbb{K}_{j-1}$) is the value of the ratio $\Psi_{2}$.
 The (maximal) dimension of the set of values of $\alpha_{\ell^{\prime}}$ is $a_{b^{*}}^{(j)}$, where $b^{*}=v-\ell^{\prime}M_{\rho_{j}}$.
-Since $1$, $z_{0}$, …$z_{0}^{\ell^{\prime}}$ are linear independent over $\mathbb{K}_{j-1}$, the total dimension is equal to $\ell^{\prime}[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b^{*}}^{(j)}$.
+Since $1$, $z_{0}$, $z_{0}^{\ell^{\prime}}$ are linear independent over $\mathbb{K}_{j-1}$, the total dimension is equal to $\ell^{\prime}[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a_{b^{*}}^{(j)}$.
 Thus it is maximal only for $\ell^{\prime}=\ell$ and is equal to $\ell[\mathbb{K}_{j-1}:\mathbb{K}_{0}]+a^{(j)}_{b}$.
 Lemma 6.
-If $\rho_{j}<\varkappa\leq\delta_{C}$ (that is $\varkappa$ is on the geodesic between $\rho_{j}$ and $\delta_{C}$ and $\varkappa\neq\rho_{j}$), then $a_{M_{\varkappa}}^{(j+1)}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$.
+If $\rho_{j}<\varkappa\leq\delta_{C}$ (that is $\varkappa$ is on the geodesic between $\rho_{j}$ and $\delta_{C}$ and $\varkappa\neq\rho_{j}$; in particular for $\varkappa=\rho_{j+1}$), then $a_{M_{\varkappa}}^{(j+1)}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$.
 Proof..
@@
 On the component $E_{\varkappa}$ this ratio is of the form $c\frac{w-\alpha}{w-1}$ with $c\in\mathbb{K}$ and (arbitrary!) $\alpha\in\mathbb{K}_{j}$.
-The value on $\delta_{C}$ (equal to the value at $w=0$, i.e.
-$c\alpha$) is an arbitrary point of $\mathbb{K}_{j}$.
+The value on $\delta_{C}$ (equal to the value at $w=0$, i.
+e., $c\alpha$) is an arbitrary point of $\mathbb{K}_{j}$.
 Therefore $d^{\{k^{0}_{\sigma}\}}=[\mathbb{K}_{j}:\mathbb{K}_{0}]$ and $a_{M_{\varkappa}}^{(j+1)}\geq[\mathbb{K}_{j}:\mathbb{K}_{0}]$.
@@
 If $v\in S^{\mathbb{K}}_{C}$ cannot be represented as $M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, one has $a_{v}^{j+1}=a_{v}^{j}$, what is compatible with (14).
-Let $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, $11\leq\ell<\ell_{j}-$ 1, and $v$ cannot be represented as $(\ell+1)M_{\rho_{j}}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$.
+Let $v=\ell M_{\rho_{j}}+b$ with $b\in S^{\mathbb{K}}_{C}$, $1\leq\ell<\ell_{j}-1$, and $v$ cannot be represented as $(\ell+1)M_{\rho_{j}}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$.
 In this case Equation (14) reduces to
@@
 (all other summands are equal to zero).
-Since, for $10\leq\ell^{\prime}\leq\ell-$ 1, $v-\ell^{\prime}M_{\rho_{j}}=M_{\rho_{j}}+b^{*}$ with $b^{*}\in S^{\mathbb{K}}_{C}$ and, due to Lemma 7, $a_{M_{\rho_{j}}}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$, all the summands in (11) except the last one are equal to $[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$.
+Since, for $0\leq\ell^{\prime}\leq\ell-1$, $v-\ell^{\prime}M_{\rho_{j}}=M_{\rho_{j}}+b^{*}$ with $b^{*}\in S^{\mathbb{K}}_{C}$ and, due to Lemma 6 (Lemma 3 for $j=1$), $a_{M_{\rho_{j}}}^{(j)}=[\mathbb{K}_{j-1}:\mathbb{K}_{0}]$, all the summands in (15) except the last one are equal to $\mathbb{K}_{j-1}:\mathbb{K}_{0}]$.
 Thus in this case (14) follows from Lemma 5.
@@
   1. 1)
-    The value $v$ can be represented as $M_{\varkappa}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$, $\varkappa>\rho_{j}$. In this case (10) follows from Lemmas 7 and  8.
+    The value $v$ can be represented as $M_{\varkappa}+b^{\prime}$ with $b^{\prime}\in S^{\mathbb{K}}_{C}$, $\varkappa>\rho_{j}$. In this case (14) follows from Lemmas 6 and  7.
   2. 2)
@@
 $\widetilde{P}^{(j)}_{C}(t)=P^{(j+1)}(t)=P^{(j)}(t)\cdot\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}.$ (16)
-Propositions 2 and 3 imply the statement of Theorem 2 in the case under consideration: $\rho_{1}\neq\sigma_{0}$.
-In the case $\rho_{1}=\sigma_{0}$, Lemmas 58 and thus Proposition 3 hold for $j\geq 2$.
+Proof of Theorem 2.
+If $\rho_{1}\neq\sigma_{0}$, the statement follows directly from Propositions 3 and 4.
+If $\rho_{1}=\sigma_{0}$, Lemmas 47 and thus Proposition 4 hold for $j\geq 2$.
 Proposition 3 does not hold.
@@
 Thus the statement of Theorem 2 holds in this case as well.
-∎
+$\Box$
 Remark 11.
 Any series $\zeta(t)=1+c_{1}t+c_{2}t^{2}+\cdots\in\mathbb{Z}[[t]]$ can be in a unique way represented as $\zeta(t)=\prod_{m=1}^{\infty}(1-t^{m})^{s_{m}}$ with $s_{m}\in\mathbb{Z}$.
-We say that the series is cyclotomic if this product/ratio of binomials is finite.
-There is a natural generalization of this notion for series in sevearl variables.
-Ussually in previous computations (say e.g.
-[5]) the Poincaré series appeared to be cyclotomic.
-This is not the case for the classical Poincaré series $P_{C}(t)$ in case II, i.e.
-when $s=\infty$.
+We say that the series is cyclotomic if this product/ratio of the binomials is finite.
+There is a natural generalization of this notion for series in several variables.
+Usually in previous computations (say, e.g., [4]) the Poincaré series appeared to be cyclotomic.
+This is not so for the Poincaré series $P^{\mathbb{K}}_{C}(t)$ in case II, i.
+e., when $s=\infty$.
 Remark 12.
-Let us assume that the curve valuation is of type I.
-Let $c$ be the conductor of the semigroup $S^{\mathbb{K}}_{C}$ (i.e the minimal element of $S^{\mathbb{K}}_{C}$ such that $c+\mathbb{Z}_{\geq 0}\subset S^{\mathbb{K}}_{C}$).
-It is known that $c$ can be expresed as
+Let us assume that the curve valuation corresponds to Case I.
+Let $c$ be the conductor of the semigroup $S^{\mathbb{K}}_{C}$ (i.e., the minimal element of $S^{\mathbb{K}}_{C}$ such that $c+\mathbb{Z}_{\geq 0}\subset S^{\mathbb{K}}_{C}$).
+It is known that $c$ can be expressed as
 $c=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1\;.$
-Let $\Delta:=c+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}$ and $\ell=\ell_{1}\cdots\ell_{s}=[\mathbb{K}_{s}:\mathbb{K}]$.
+Let $\ell=\ell_{1}\cdots\ell_{s}=[\mathbb{K}_{s}:\mathbb{K}]$ and
+$\Delta:=c+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}\;.$
 The computation of the Poincaré series gives that $a_{v}=\dim J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)=\ell$ for all $v\geq\Delta$ and $a_{\Delta-1}<\ell$.
-Moreover, one has the following symmetry property for all $v\in\mathbb{Z}$:
-$\dim\left({J^{\mathbb{K}}(v)}\left/{J^{\mathbb{K}}(v+1)}\right.\right)+\dim\left({J^{\mathbb{K}}(\Delta-1-v)}\left/{J^{\mathbb{K}}(\Delta-v)}\right.\right)=\ell\;$ (13)
-Following the notations of Remark 2, the integral closure $\overline{R}$ of the local ring $R=\mathbb{K}[[x,y]]/(F)$ of the algebroid curve defined by $F$ is a valuation ring and is isomophic to $\mathbb{K}^{\prime}[[u]]$ ($u$ is an uniformizing parameter).
+Moreover, one has the following symmetry property:
+Corollary 1.
+For all $v\in\mathbb{Z}$ one has
+[…]
+Following the notations of Remark 4, the integral closure $\overline{R}$ of the local ring $R=\mathbb{K}[[x,y]]/(F)$ of the algebroid curve defined by $F$ is a valuation ring and is isomorphic to $\mathbb{K}^{\prime}[[u]]$ ($u$ is an uniformizing parameter).
 One has:
 Corollary 2.
-Let
-$\Delta=\sum_{i=1}^{g}(N_{i}-1)M_{\sigma_{i}}-M_{\sigma_{0}}+1+\sum_{j=1}^{s}(\ell_{j}-1)M_{\rho_{j}}\;.$
-Then, the conductor ideal ${\mathcal{C}}(\overline{R}/R)=\{z\in\overline{R}:z\overline{R}\subset R\}$ is the ideal $u^{\Delta}\overline{R}$.
-Note that the ring $R$ is Gorenstein (being plane) and the symmetry of the equation (13) is the just the characterization of the Gorenstein condition of the ring $R$ (cf. [3]).
+The conductor ideal ${\mathcal{C}}(\overline{R}/R)=\{z\in\overline{R}:z\overline{R}\subset R\}$ is the ideal $u^{\Delta}\overline{R}$.
+Note that the ring $R$ is Gorenstein (being plane) and the symmetry described by the equation (17) is just the characterization of the Gorenstein condition of the ring $R$ (cf. [3]).
 7 The Poincaré series of a divisorial valuation
 Let $\nu=\nu_{\delta}$ be a divisorial valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$.
-This means that it is defined by a component $E_{\delta}$ of the exceptional divisor of a modification of $(\mathbb{C}^{2},o)$.
-Let $\pi:(W,{\mathcal{D}})\to(\mathbb{C}^{2},0)$ be the minimal $\mathbb{K}$-resolution of the valuation $\nu$.
+This means that it is defined by a component $E_{\delta}$ of the exceptional divisor of a modification of $(\mathbb{C}^{2},0)$.
+We assume that the corresponding $G$-invariant modification is finite.
+Otherwise we are in a situation similar to Case III in the curve case and the valuation $\nu_{\delta}$ is a multiple of another one: corresponding to the component on which the $G$-orbit of the blown-up point is infinite.
+Let $\pi:(W,{\mathcal{D}})\to(\mathbb{C}^{2},0)$ be the minimal $G$-resolution of the valuation $\nu$.
 The quotient of the dual graph of this resolution by the action of the Galois group $G$ looks like in figure 7.
@@
 This graph essentially coincides with the minimal resolution graph of a $\mathbb{K}_{\delta}$-curvette at the component $E_{\delta}$ ($\mathbb{K}_{\delta}=\mathbb{K}_{s}$) with a possible tail from $\tau_{g}$ to $\delta$ added.
-Here, as above, $\sigma_{i}$, $i=0,1,\ldots,g$ are the dead ends of the graph $\check{\Gamma}$, $\tau_{i}$, $i=1,\ldots,g$ are the rupture points and $\rho_{j}$, $j=1,\ldots,s$ are the branching points of the $\mathbb{K}$-resolution.
+Here, as above, $\sigma_{i}$, $i=0,1,\ldots,g$, are the dead ends of the graph $\check{\Gamma}$, $\tau_{i}$, $i=1,\ldots,g$, are the rupture points and $\rho_{j}$, $j=1,\ldots,s$, are the splitting points of the $G$-resolution.
 Remark 13.
@@
 We shall keep the notations ${\mathcal{E}}^{\{k_{\sigma}\}}_{\mathbb{K},0}$, $F^{\{k_{\sigma}\}}$, and $d^{\{k_{\sigma}\}}$ as before.
+Let $C$ be a $\mathbb{K}_{s}$-curvette at the component $E_{\delta}$.
 Theorem 3.
 One has
-$P_{\nu}(t)=P^{S}(t)\frac{1}{1-t^{M_{\delta}}}\cdot\prod_{i=1}^{s}\frac{1-t^{\ell_{i}M_{\rho_{i}}}}{1-t^{M_{\rho_{i}}}}\;.$ (14)
+$P^{\mathbb{K}}_{\nu_{\delta}}(t)=P^{S}_{C}(t)\cdot\frac{1}{1-t^{M_{\delta}}}\cdot\prod_{i=1}^{s}\frac{1-t^{\ell_{i}M_{\rho_{i}}}}{1-t^{M_{\rho_{i}}}}\;.$ (18)
 Proof..
-Let $C$ be a $\mathbb{K}_{s}$-curvette at the component $E_{\delta}$.
-For the coefficient $a_{v}$ in the Poincaré series $P_{\nu}(t)=\sum a_{v}t^{v}$ one has $a_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}$ such that $\nu(\{k_{\sigma}\})=\sum k_{\sigma}M_{\sigma}=v$.
-Let $A^{\prime}$ be the set of collections $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$ such that $k_{\delta}=0$ and let $P^{\prime}(t)=\sum a^{\prime}_{v}t^{v}$ with $a^{\prime}_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximun is taken over the collections $\{k_{\sigma}\}\in A^{\prime}$ such that $\nu(\{k_{\sigma}\})=v$.
+For the coefficient $a_{v}$ in the Poincaré series $P^{\mathbb{K}}_{\nu_{\delta}}(t)=\sum a_{v}t^{v}$ one has $a_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximum is taken over the collections $\{k_{\sigma}\}$ such that $\nu(\{k_{\sigma}\})=\sum k_{\sigma}M_{\sigma}=v$.
+Let $A^{\prime}$ be the set of collections $\{k_{\sigma}\}$, $\sigma\in\check{\Gamma}$ such that $k_{\delta}=0$ and let $P^{\prime}(t)=\sum a^{\prime}_{v}t^{v}$ with $a^{\prime}_{v}:=\max d^{\{k_{\sigma}\}}$ where the maximum is taken over the collections $\{k_{\sigma}\}\in A^{\prime}$ such that $\nu(\{k_{\sigma}\})=v$.
 Lemma 8.
 One has
-$P^{\prime}_{\nu}(t)=P_{C}(t)=P^{S}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}\;.$
+$P^{\prime}_{\nu}(t)=P_{C}^{\mathbb{K}}(t)=P^{S}_{C}(t)\cdot\prod_{j=1}^{s}\frac{1-t^{\ell_{j}M_{\rho_{j}}}}{1-t^{M_{\rho_{j}}}}\;.$
 Proof..
@@
 Lemma 9.
-Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\delta}+b$ with $b\in S^{\mathbb{K}}_{C}$ and the maximal possible $\ell$ in this representation.
+Let $v\in S^{\mathbb{K}}_{C}$ be of the form $v=\ell M_{\delta}+b$ with $b\in S^{\mathbb{K}}_{C}$ and such that $b-M_{\delta}\notin S_{C}^{\mathbb{K}}$.
 Then $a_{v}=\ell[\mathbb{K}_{s}:\mathbb{K}_{0}]+a^{\prime}_{b}$.
 Proof..
-(Cf.
-proof of Lemma 6.) Let us represent a collection $\{k_{\sigma}\}$ with $\nu(\{k_{\sigma}\})=v$ in the form $k_{\sigma}=k^{1}_{\sigma}+k^{2}_{\sigma}$, where $k^{1}_{\delta}=\ell^{\prime}$ ($\ell^{\prime}\leq\ell$), $k^{1}_{\sigma}=0$ for $\sigma\neq\delta$, $k^{2}_{\delta}=0$.
-A function $f\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k_{\sigma}\}}$ can be represented as $f=f_{1}f_{2}$ with $f_{r}\in{\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $r=1,2$.
-Let $h_{r}$, $r=1,2$, be fixed functions from ${\mathcal{E}}_{\mathbb{K}^{2},0}^{\{k^{r}_{\sigma}\}}$, $h=h_{1}h_{2}$.
-As in Lemma 6, we shall use a coordinate on $E_{\delta}$ such that the intersection point with the previous component from the geodesic to $\sigma_{0}$ is equal to $\infty$ and we shall take the function $h_{1}$ described there (that is, the $\ell^{\prime}$th power of the left hand side of the equation of the $G_{s}$-curvette at the point $w=1$ of the component $E_{\delta}$).
-One has $\Psi={\widetilde{f}}/{\widetilde{h}}=\Psi_{1}\Psi_{2}$, where $\Psi_{r}={\widetilde{f}_{r}}/{\widetilde{h}_{r}}$.
-The function $\Psi_{2}$ is constant on the component $E_{\delta}$.
-The maximal possible dimension (over $\mathbb{K}$) of the set of its values on this component is $a^{\prime}_{b^{\prime}}$ ($b^{\prime}=v-\ell^{\prime}M_{\delta}$).
-The function $\Psi_{1}$ is constant on the union of all the components $E_{\sigma}$ for $\sigma$ from the geodesic between $\sigma_{0}$ and $\delta$, excluding the last one.
-Its restriction to the component $E_{\delta}$ is equal to $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})/(w-1)^{\ell^{\prime}}$ with $c\in\mathbb{K}$, where $w_{p}\in\mathbb{K}_{s}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\delta}$.
+[…]
+$\frac{c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})}{(w-1)^{\ell^{\prime}}}$
+with $c\in\mathbb{K}$, where $w_{p}$ are the intersection points of the strict transform of the curve $\{f_{1}=0\}$ with the component $E_{\delta}$.
+The set of these points is an arbitrary union of $G_{s}$-orbits.
 Therefore, up to the fixed multiple $(w-1)^{\ell^{\prime}}$, the restriction of $\Psi$ to the component $E_{\delta}$ is equal to the restriction of $\Psi_{2}$ multiplied by $c\prod_{p=1}^{\ell^{\prime}}(w-w_{p})$.
-The dimension (over $\mathbb{K}$) of the space of functions of the latter form is $\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}]$ (since $w_{p}$ are arbitrary elements of $\mathbb{K}_{s}$).
+The dimension (over $\mathbb{K}$) of the space of functions of the latter form is $\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}]$ (since the collection of the symmetric polynomials of the points $w_{p}$ is a generic point of $\mathbb{K}_{s}^{\ell^{\prime}}$).
 Therefore the (maximal) dimension of the set of these functions on $E_{\delta}$ is $a^{\prime}_{b^{\prime}}+\ell^{\prime}[\mathbb{K}_{s}:\mathbb{K}_{0}]$.
@@
 One has
-$P_{\nu}(t)=P^{\prime}_{\nu}(t)\cdot\frac{1}{1-t^{M_{\delta}}}.$
+$P^{\mathbb{K}}_{\nu_{\delta}}(t)=P^{\prime}(t)\cdot\frac{1}{1-t^{M_{\delta}}}=P^{\mathbb{K}}_{C}(t)\cdot\frac{1}{1-t^{M_{\delta}}}.$
 Lemma 8 and Proposition 5 imply the statement of Theorem 3. ∎
 Remark 14.
-One can see that the classical Poincaré series of a curve or of a divisorial valuation determines the combinatorial type of the minimal $\mathbb{K}$-resolution of the valuation.
+One can see that the Poincaré series of a curve or of a divisorial valuation determines the combinatorial type of the minimal $G$-resolution of the valuation.
+Remark 15.
+The formula in Proposition 5 relating the Poincaré series of the divisorial valuation $\nu$ and the Poincaré series of a $\mathbb{K}_{s}$-curvette $C$ at the component $E_{\delta}$ obviously extends the known one for the complex case (see [12] or [10]).
+Even the formula itself and, to some extent, the proofs have some technical resemblance with the relation of both series in the case of several divisorial valuations in [9].
+However, the objects are rather diferent, in [9] all the considerations are in the pure complex analytic framework and the proofs use only properties of pencils (another way to treat meromorphic functions like in this paper) in the complex setting.